### v1

# Multichannel Conflict-Avoiding Codes for Expanded Scenarios

Abstract
A conflict-avoiding code (CAC) of length $L$ and weight $w$ is used for deterministic multiple-access without feedback. When the number of simultaneous active users is less than or equal to $w$, such a code is able to provide a hard guarantee that each active user has a successful transmission within every consecutive $L$ time slots. Recently, CACs were extended to multichannel CAcs (MC-CACs) over $M>1$ orthogonal channels with the aim of increasing the number of potential users that can be supported. While most existing results on MC-CAC are derived under the assumption $M\geq w$, this paper focuses on the case $M<w$, which is more relevant to practical application scenarios. In this paper, we first introduce the concept of exceptional codewords in MC-CACs. By employing Kneser’s Theorem and some techniques from additive combinatorics, we derive a series of optimal MC-CACs. Along the way, several previously known optimal CAC results are generalized. Finally, our results extend naturally to AM-OPPTS MC-CACs and mixed-weight MC-CACs, two classes of relevant codes.
I Introduction
I-A Single-Channel Conflict-Avoiding Codes
Let $\mathbb{Z}_{L}\triangleq\{0,1,\ldots,L-1\}$ denote the ring of residue modulo $L$. For $S\subseteq\mathbb{Z}_{L}$, let
$d^{*}(S)\triangleq\{a-b\,(\text{mod }L):\,a,b\in S,a\neq b\}$
denote the set of (nonzero) differences of $S$.
Definition 1.
Let $L$ and $w$ be two positive integers with $L\geq w$. A conflict-avoiding code (CAC) $\mathcal{C}$ of length $L$ with weight $w$ is a collection of $w$-subsets, called codewords, of $\mathbb{Z}_{L}$ such that
$d^{*}(S)\cap d^{*}(S^{\prime})=\emptyset\quad\forall S,S^{\prime}\in\mathcal{C},S\neq S^{\prime}.$ (1)
The equation (1) is called the disjoint-difference-set property.
Let $\mathrm{CAC}(L,w)$ denote the class of all CACs of length $L$ with weight $w$. The maximum size of a code in $\mathrm{CAC}(L,w)$ is denoted by $K(L,w)$, i.e.,
$K(L,w)\triangleq\max\{|\mathcal{C}|:\,\mathcal{C}\in\mathrm{CAC}(L,w)\}.$
A code $\mathcal{C}\in\mathrm{CAC}(L,w)$ is called optimal if its code size achieves $K(L,w)$.
A $w$-subset $S\subseteq\mathbb{Z}_{L}$ is said to be equi-difference with generator $g\in\mathbb{Z}_{L}^{*}$ if $S$ is of the form $\{0,g,2g,\ldots,(w-1)g\}$. A CAC is called equi-difference if it entirely consists of equi-difference codewords. Let $\mathrm{CAC^{e}}(L,w)\subset\mathrm{CAC}(L,w)$ denote the class of all equi-difference codes and $K^{e}(L,w)$ be the maximum size among $\mathrm{CAC^{e}}(L,w)$. Obviously, $K^{e}(L,w)\leq K(L,w)$.
I-B Multichannel Conflict-Avoiding Codes
In a multiple access network with $M$ channels, a transmission pattern is represented as a zero-one $M\times L$ array, where $M$ denotes the number of channels and $L$ refers to the codeword length. The rows and columns of the array are further indexed by $\mathcal{I}_{M}\triangleq\{1,2,...,M\}$ and $\mathbb{Z}_{L}$, respectively. In this setting, a codeword with weight $w$ can be realized as a $w$-subset in $\mathcal{I}_{M}\times\mathbb{Z}_{L}$, in which each element is an ordered pair.
For two indices $i,j\in\mathcal{I}_{M}$, define the $(i,j)$ set of differences by
$D_{S}(i,j):=\begin{cases}\{t_{1}-t_{2}\ (\text{mod }L):\,t_{1}\in S_{i},t_{2}\in S_{j}\}&\text{if }i\neq j,\\ \{t_{1}-t_{2}\ (\text{mod }L):\,t_{1}\in S_{i},t_{2}\in S_{j}\}\setminus\{0\}&\text{if }i=j.\end{cases}$ (2)
Note that, in the identical-index case, $D_{S}(i,i)=d^{*}(S_{i})$. $D_{S}$ is called the array of differences of the transmission pattern $S$. Two arrays of differences $D_{S}$ and $D_{S^{\prime}}$ are said to be disjoint, denoted by $D_{S}\cap D_{S^{\prime}}=\emptyset,$ if they are entry-wise disjoint.
Definition 2.
Let $M$, $L$ and $w$ be positive integers. A multichannel conflict-avoiding code (MC-CAC) $\mathcal{C}$ of length $L$ with weight $w$ in $M$ channels is a collection of $w$-subsets, also called codewords, of $\mathcal{I}_{M}\times\mathbb{Z}_{L}$ such that
$D_{S}\cap D_{S^{\prime}}=\emptyset\quad\forall S,S^{\prime}\in\mathcal{C},S\neq S^{\prime}.$ (3)
The equation (3) is called the disjoint-difference-array property.
Let $\text{MC-CAC}(M,L,w)$ denote the class of all MC-CACs of length $L$ with weight $w$ in $M$ channels. The maximum size of a code in $\text{MC-CAC}(M,L,w)$ is denoted by $K(M,L,w)$, i.e.,
$K(M,L,w)\triangleq\max\{|\mathcal{C}|:\,\mathcal{C}\in\text{MC-CAC}(M,L,w)\}.$
A code $\mathcal{C}\in\text{MC-CAC}(M,L,w)$ is called optimal if its code size reaches $K(M,L,w)$.
I-C Contribution
The technical results and main contributions of this paper are summarized below.
  1. 1.
    Define the so-called exceptional codewords in MC-CACs, and derive a series of related properties by means of Kneser’s theorem. See Lemma 2 and Lemma 3 in Section II.
  2. 2.
    Obtain two classes of optimal single-channel CACs of lengths $p_{1}^{r_{1}}\cdots p_{n}^{r_{n}}$ and $(w-1)p_{1}^{r_{1}}\cdots p_{n}^{r_{n}}$, where $p_{i}$ are primes and $w$ refers to the Hamming weight. See Theorems 7 and 10 in Section IV. The constructions of these codes can be extended to mixed-weight CACs.
  3. 3.
    Obtain optimal MC-CACs of length $(w-1)p_{1}^{r_{1}}\cdots p_{n}^{r_{n}}$ with two channels. See Theorem 13 in Section V. The construction of these codes can be extended to mixed-weight MC-CACs, which are defined for the first time in this paper.
  4. 4.
    Propose a construction of MC-CACs of length $(2\frac{w}{M}-1)L^{\prime}$, where $w$ and $M$ refer to the Hamming weight and the number of channels. See Theorem 14 in Section VI.
  5. 5.
    Derive a general upper bound on the maximum size of an MC-CAC, where the code length is with some restriction. This can be seen as a generalization of some previously-known results. See Theorem 4 in Section III.
  6. 6.
    Some applications to MC-CACs with at most one-packet per time slot property. See Theorems 17, 18 and 19 in Section VII-A.
II Exceptional Codewords
In a (single-channel) CAC in $\mathrm{CAC}(L,w)$, a codeword $S$ is called exceptional if $|d^{*}(S)|<2w-2$ [1, 2, 5]. In this section, we will extend the concept of exceptional codewords to MC-CACs, together with some characterizations of structural properties.
II-A Exceptional Pairs and Revisit of the Kenser’s Theorem
We first give some notion of Additive Combinatorics [6]. For two subsets $A,B\subseteq\mathbb{Z}_{L}$ and an element $x\in\mathbb{Z}_{L}$, define
$x+A$ $\triangleq\{x+a:\,a\in A\},$
$A+B$ $\triangleq\{a+b:\,a\in A,b\in B\},\text{ and}$ (4)
$A-B$ $\triangleq\{a-b:\,a\in A,b\in B\}.$
Moreover, define
$d(A)\triangleq A-A.$
Note that $0\in d(A)$ and $d(A)\setminus\{0\}=d^{*}(A)$, the set of nonzero differences of $A$. In this notion, the definition of exceptional codewords $S$ in a CAC can be represented by $|S-S|\leq 2|S|-2$. We first extend this concept to a pair of subsets.
Definition 3.
A pair of two subsets $\{A,B\}$ is called exceptional if
$|A-B|\leq|A|+|B|-2.$ (5)
Note that |AB|=|BA||A-B|=|B-A|, so there is no confusion in writing ABA-B or BAB-A in (5).
Let $T$ be a non-empty subset in $\mathbb{Z}_{L}$. The set of stabilizers of $T$ in $\mathbb{Z}_{L}$ is defined as
$\mathsf{H}(T)\triangleq\{h\in\mathbb{Z}_{L}:\,h+T=T\}.$
It is obvious that $0\in\mathsf{H}(T)$ and $\mathsf{H}(T)$ is a subgroup of $\mathbb{Z}_{L}$. So, $|\mathsf{H}(T)|$ divides $L$ by Lagrange’s theorem. $T$ is called periodic if $\mathsf{H}(T)$ is non-trivial, namely, $\mathsf{H}(T)\neq\{0\}$. Here we list some well-known properties (e.g., see [2, 5]) of the set of stabilizers.
Proposition 1.
Let $T\subseteq\mathbb{Z}_{L}$ be non-empty.
  1. (i)
    $\mathsf{H}(T)$ is a subgroup of $\mathbb{Z}_{L}$, and thus $|\mathsf{H}(T)|$ divides $L$.
  2. (ii)
    If $0\in T$, then $\mathsf{H}(T)\subseteq T$.
  3. (iii)
    If $T$ is periodic, then $T=\bigcup_{a\in T}(a+\mathsf{H}(T))$. Moreover, $|\mathsf{H}(T)|$ divides $|T|$.
The following lists some results about the number of stabilizers of the difference set of an exceptional subset, which can be seen in [5, Corollary 1 and Lemma 1].
Lemma 1 ([5]).
Let $A$ be a nonempty subset in $\mathbb{Z}_{L}$. If $A$ is exceptional, then
  1. (i)
    $2\leq|\mathsf{H}(A-A)|\leq 2|A|-2$,
  2. (ii)
    $|\mathsf{H}(A-A)|$ does not divide 2|A|12|A|-1, and
  3. (iii)
    $|\mathsf{H}(A-A)|$ does not divide |A|1|A|-1.
The above results were based on Kneser’s theorem, in which the formal description is given below.
Theorem 1 ([7, 6]).
Let $A$ and $B$ be two non-empty subsets in $\mathbb{Z}_{L}$, and let $\mathsf{H}=\mathsf{H}(A+B)$. Then,
$|A+B|\geq|A+\mathsf{H}|+|B+\mathsf{H}|-|\mathsf{H}|.$ (6)
In particular,
$|A+B|\geq|A|+|B|-|\mathsf{H}|.$ (7)
We shall study the number of stabilizers of the difference set of an exceptional pair, which can be seen as generalizations of Lemma 1.
Lemma 2.
Let $A,B$ be two nonempty subsets of $\mathbb{Z}_{L}$. If $\{A,B\}$ is an exceptional pair, then
  1. (i)
    $2\leq|\mathsf{H}(A-B)|\leq|A|+|B|-2$, and
  2. (ii)
    $|\mathsf{H}(A-B)|$ does not divide |A|+|B|1|A|+|B|-1.
  3. (iii)
    $|\mathsf{H}(A-B)|$ does not divide |A|1|A|-1 when $|A|=|B|$.
Proof.
For notational convenience, denote by $\mathsf{H}=\mathsf{H}(A-B)$.
(i) By (5) and (7), we have
$|A|+|B|-2$ $\geq|A-B|=|A+(-B)|$
$\geq|A|+|-B|-|\mathsf{H}|=|A|+|B|-|\mathsf{H}|,$
which implies that $|\mathsf{H}|\geq 2$. Therefore, the set ABA-B is periodic. By Proposition 1(iii), we get $|\mathsf{H}|\leq|A-B|$, which is less than or equal to |A|+|B|2|A|+|B|-2 because $\{A,B\}$ is exceptional.
(ii) Suppose to the contrary that $|\mathsf{H}|$ divides |A|+|B|1|A|+|B|-1. Since, by Proposition 1(i), $\mathsf{H}$ is a subgroup of $\mathbb{Z}_{L}$, we have $\mathsf{H}=-\mathsf{H}$, which implies that $|-B+\mathsf{H}|=|-(B+\mathsf{H})|=|B+\mathsf{H}|$. Moreover, as $B+\mathsf{H}$ is the disjoint union of cosets of $\mathsf{H}$, the size $|B+\mathsf{H}|$ is a multiple of $|\mathsf{H}|$, and thus $|B+\mathsf{H}|\geq|\mathsf{H}|\cdot\left\lceil|B|/|\mathsf{H}|\right\rceil$. Similarly, $|A+\mathsf{H}|\geq|\mathsf{H}|\cdot\left\lceil|A|/|\mathsf{H}|\right\rceil$. By plugging $A=A$, B=BB=-B into (6) yields that
$|A-B|\geq|A+\mathsf{H}|+|B+\mathsf{H}|-|\mathsf{H}|\geq|\mathsf{H}|\left(\left\lceil\frac{|A|}{|\mathsf{H}|}\right\rceil+\left\lceil\frac{|B|}{|\mathsf{H}|}\right\rceil-1\right).$
Since $\{A,B\}$ is an exceptional pair, namely, $|A-B|\leq|A|+|B|-2$, it follows that
$|A|+|B|-2\geq|\mathsf{H}|\left(\left\lceil\frac{|A|}{|\mathsf{H}|}\right\rceil+\left\lceil\frac{|B|}{|\mathsf{H}|}\right\rceil-1\right).$ (8)
Let $|A|=q_{1}|\mathsf{H}|+r_{1}$ and $|B|=q_{2}|\mathsf{H}|+r_{2}$, for some integers $q_{1},q_{2}\geq 0$ and $0<r_{1},r_{2}\leq|\mathsf{H}|$. Note that
$1<r_{1}+r_{2}\leq 2|\mathsf{H}|.$ (9)
By (8), we have
$(q_{1}+q_{2})|\mathsf{H}|+r_{1}+r_{2}-2=|A|+|B|-2\geq|\mathsf{H}|\big((q_{1}+1)+(q_{2}+1)-1\big),$
which implies that
$r_{1}+r_{2}-2\geq|\mathsf{H}|.$ (10)
Combining (9) and (10) yields
$|\mathsf{H}|+1\leq r_{1}+r_{2}-1\leq 2|\mathsf{H}|-1.$ (11)
By the assumption that $|\mathsf{H}|$ divides $|A|+|B|-1=|\mathsf{H}|(q_{1}+q_{2})+r_{1}+r_{2}-1$, we get $|\mathsf{H}|$ divides $r_{1}+r_{2}-1$. This contradicts to (11).
(iii) This case can be dealt with in the same argument we provided for (ii). By letting $|A|=|B|=q|\mathsf{H}|+r$, i.e., $q=q_{1}=q_{2}$ and $r=r_{1}=r_{2}$, the two inequalities (9) and (10) imply
$\frac{|\mathsf{H}|}{2}\leq r-1\leq|\mathsf{H}|-1.$ (12)
Suppose to the contrary that $|\mathsf{H}|$ divides $|A|-1=q|\mathsf{H}|+r-1$. We get $|\mathsf{H}|$ divides r1r-1, which is a contradiction to (12). Hence the proof is completed. ∎
II-B Exceptional Codewords in MC-CACs
Consider a subset $S\in\mathcal{I}_{M}\times\mathbb{Z}_{L}$. For $i\in\mathcal{I}_{M}$, define
$S_{i}\triangleq\{t:\,(i,t)\in S\}.$
Obviously, $|S|=\sum_{i\in\mathcal{I}_{M}}|S_{i}|$. We further define
$e_{S}\triangleq|\{i\in\mathcal{I}_{M}:\,S_{i}\neq\emptyset\}|$
indicating the number of channels that have at least one packet in the transmission pattern $S$. With the notation given in (4), the $(i,j)$ set of difference of $S$, $D_{S}(i,j)$, defined in (2) can be rewritten as $S_{i}-S_{j}$ if $i\neq j$.
We are ready to define exceptional subsets in $\mathcal{I}_{M}\times\mathbb{Z}_{L}$.
Definition 4.
A subset $S\in\mathcal{I}_{M}\times\mathbb{Z}_{L}$ is called exceptional if at least one of the followings holds.
  1. (i)
    There exists $i\in\mathcal{I}_{M}$ such that $S_{i}$ is exceptional, i.e., $|S_{i}-S_{i}|\leq 2|S_{i}|-2$.
  2. (ii)
    There exist distinct indices $i,j\in\mathcal{I}_{M}$ such that $\{S_{i},S_{j}\}$ is exceptional, i.e., $|S_{i}-S_{j}|\leq|S_{i}|+|S_{j}|-2$.
Lemma 3.
Let $S$ be a $w$-subset in $\mathcal{I}_{M}\times\mathbb{Z}_{L}$. If the prime factors of $L$ are all larger than or equal to 2w12w-1, then $S$ is not exceptional.
Proof.
Suppose to the contrary that $S$ is exceptional. Let $T$ be any nonempty subset in $\mathbb{Z}_{L}$. Since all prime factors $L$ are all larger than or equal to 2w12w-1, by Proposition 1(i), we have
$|\mathsf{H}(T)|=1\quad\text{or}\quad|\mathsf{H}(T)|\geq 2w-1.$ (13)
If there exists an index $i\in\mathcal{I}_{M}$ such that $S_{i}$ is exceptional, by Lemma 1(i), we have $2\leq|\mathsf{H}(S_{i}-S_{i})|\leq 2|S_{i}|-2\leq 2w-2$. If there exist distinct indices $i,j\in\mathcal{I}_{M}$ such that $\{S_{i},S_{j}\}$ is exceptional, by Lemma 2(i), we have $2\leq|\mathsf{H}(S_{i}-S_{j})|\leq|S_{i}|+|S_{j}|-2\leq w-2$. Either case contradicts to (13), so the proof is done. ∎
We end this section with a fundamental result in Group Theory, which will be used in subsequent sections.
Proposition 2.
The subgroup of $\mathbb{Z}_{L}$ is uniquely determined by its order. More precisely, for any divisor $d$ of $L$, the unique subgroup of $\mathbb{Z}_{L}$ with order $d$ is $\{0,L/d,2L/d,\ldots,(d-1)L/d\}$.
III General Upper Bounds on the Maximum Code Size
We first recall an upper bound on the maximum code size of a CAC with some restriction on the code length.
Theorem 2 ([1], Theorem 5).
Let $L$ and $w$ be positive integers. If the prime factors of $L$ are all larger than or equal to 2w12w-1, then
$K(L,w)\leq\frac{L-1}{2w-2}.$
The following result is due to [4].
Theorem 3 ([4], Corollary 13).
Let $M,L$ and $w$ be positive integers. If the prime factors of $L$ are all larger than or equal to 2w12w-1, we have
$K(M,L,w)\leq\frac{M(M-1)L}{w(w-1)}+\frac{M(L-1)}{2w-2}.$ (14)
The upper bounds in Theorem 3 can be improved in the case when $M<w$, which is shown in the following theorem.
Theorem 4.
Let $M,L$ and $w$ be positive integers with $M<w$. If the prime factors of $L$ are all larger than or equal to 2w12w-1, we have
$K(M,L,w)\leq\frac{M(M-1)L}{(2w-M)(w-1)}+\frac{M(L-1)}{2w-2}.$ (15)
Proof.
Let $\mathcal{C}$ be any code in $\text{MC-CAC}(M,L,w)$. The codewords in $\mathcal{C}$ are partitioned, according to the number of channels that are occupied in their transmission patterns, i.e., $e_{S}$, into the following three classes:
$\mathcal{C}_{1}=\{S\in\mathcal{C}:e_{S}=1\},$
$\mathcal{C}_{2}=\{S\in\mathcal{C}:e_{S}=M\},\text{ and}$
$\mathcal{C}_{3}=\{S\in\mathcal{C}:2\leq e_{S}\leq M-1\}.$
Since the prime factors of $L$ are all larger than or equal to 2w12w-1, by Lemma 3, each codeword in $\mathcal{C}$ is non-exceptional. Then, we have $|S_{i}-S_{j}|\geq|S_{i}|+|S_{j}|-1$ for any $i,j\in\mathcal{I}_{M}$, and in particular, $|d^{*}(S_{i})|=|S_{i}-S_{i}|-1\geq 2(|S_{i}|-1)$.
For any $i\in\mathcal{I}_{M}$, by the disjoint-difference-array property, we have
$L-1=|\mathbb{Z}^{*}_{L}|\geq\sum_{S\in\mathcal{C}}|D_{S}(i,i)|.$
Hence,
$M(L-1)$ $\geq\sum_{i\in\mathcal{I}_{M}}\sum_{S\in\mathcal{C}}|D_{S}(i,i)|$
$=\sum_{S\in\mathcal{C}}\sum_{i\in\mathcal{I}_{M},S_{i}\neq\emptyset}|d^{*}(S_{i})|\geq\sum_{S\in\mathcal{C}}\sum_{i\in\mathcal{I}_{M},S_{i}\neq\emptyset}2(|S_{i}|-1)$
$=2(w-1)|\mathcal{C}_{1}|+2(w-M)|\mathcal{C}_{2}|+2\sum_{S\in\mathcal{C}_{3}}(w-e_{S}),$
which implies that
$(w-1)|\mathcal{C}_{1}|+(w-M)|\mathcal{C}_{2}|+\sum_{S\in\mathcal{C}_{3}}(w-e_{S})\leq\frac{M(L-1)}{2}.$ (16)
Now we consider $D_{S}(i,j)$ for distinct indices $i,j\in\mathcal{I}_{M}$, by going through all codewords in $\mathcal{C}$. The sets of differences $D_{S}(i,j)$ for $S\in\mathcal{C}$ are mutually disjoint, and their union is a subset of $\mathbb{Z}_{L}$. As $D_{S}(i,j)=\emptyset$ for $S\in\mathcal{C}_{1}$, we have
$L=|\mathbb{Z}_{L}|\geq\sum_{S\in\mathcal{C}_{2}\cup\mathcal{C}_{3}}|D_{S}(i,j)|.$ (17)
Therefore,
$M(M-1)L$ $\geq\sum_{i,j\in\mathcal{I}_{M},i\neq j}\sum_{S\in\mathcal{C}_{2}\cup\mathcal{C}_{3}}|D_{S}(i,j)|$
$\geq\sum_{S\in\mathcal{C}_{2}\cup\mathcal{C}_{3}}\sum_{\begin{subarray}{c}i,j\in\mathcal{I}_{M},i\neq j\\ S_{i},S_{j}\neq\emptyset\end{subarray}}(|S_{i}|+|S_{j}|-1)$
$=\sum_{S\in\mathcal{C}_{2}\cup\mathcal{C}_{3}}\left(2(e_{S}-1)\left(\sum_{i\in\mathcal{I}_{M},S_{i}\neq\emptyset}|S_{i}|\right)-e_{S}(e_{S}-1)\right)$
$=\sum_{S\in\mathcal{C}_{2}\cup\mathcal{C}_{3}}(e_{S}-1)(2w-e_{S})$
$\geq(M-1)(2w-M)|\mathcal{C}_{2}|+\sum_{S\in\mathcal{C}_{3}}(e_{S}-1)(2w-e_{S}),$
yields that
$(M-1)(2w-M)|\mathcal{C}_{2}|+\sum_{S\in\mathcal{C}_{3}}(e_{S}-1)(2w-e_{S})\leq M(M-1)L.$ (18)
We multiply the inequality in (16) by (2wM)(2w-M) and combine the result with (18). We then get
$\begin{split}&(w-1)(2w-M)\left(|\mathcal{C}_{1}|+|\mathcal{C}_{2}|\right)+\sum_{S\in\mathcal{C}_{3}}\left(-e_{S}^{2}+(M+1)e_{S}+2w^{2}-Mw-2w\right)\\ \leq&\ \frac{M(L-1)}{2}(2w-M)+M(M-1)L.\end{split}$ (19)
When $M=2$, we have $\mathcal{C}_{3}=\emptyset$, and thus it follows from (19) that
$|\mathcal{C}|=|\mathcal{C}_{1}|+|\mathcal{C}_{2}|\leq\frac{M(M-1)L}{(2w-M)(w-1)}+\frac{M(L-1)}{2w-2},$
as desired. So, we assume $M\geq 3$ in what follows.
Define a function $f(x)=-x^{2}+(M+1)x+2w^{2}-Mw-2w$. The derivative of $f(x)$ is 2x+M+1-2x+M+1, implying that $f(x)$ is a concave function with a unique maximum at $x=\frac{M+1}{2}$. Let us focus on the interval [2,M1][2,M-1]. Since $\frac{M+1}{2}\in[2,M-1]$ for $M\geq 3$ and f(2)=(2wM)(w1)+M2=f(M1)f(2)=(2w-M)(w-1)+M-2=f(M-1), we have
$f(x)$ $\geq\min\left\{f(2),f(M-1)\right\}$
$=(2w-M)(w-1)+M-2,$
for $2\leq x\leq M-1$. Therefore,
$f(e_{S})\geq(2w-M)(w-1)+M-2\geq(2w-M)(w-1)$ (20)
since $2\leq e_{S}\leq M-1$ and $M\geq 3$ for $S\in\mathcal{C}_{3}$. Therefore, the left-hand-side of (19) turns to
$\quad\ (w-1)(2w-M)\left(|\mathcal{C}_{1}|+|\mathcal{C}_{2}|\right)+\sum_{S\in\mathcal{C}_{3}}f(e_{S})$
$\geq(w-1)(2w-M)\left(|\mathcal{C}_{1}|+|\mathcal{C}_{2}|+|\mathcal{C}_{3}|\right)$ (21)
$=(w-1)(2w-M)|\mathcal{C}|.$
Then the inequality in (15) follows. ∎
Remark 1.
If $M\geq 3$, the equality (15) will hold only when $\mathcal{C}_{3}=\emptyset$, that is, all involved codewords $S$ has either $e_{S}=1$ or $e_{S}=M$. This fact can be seen from (20) – (21): If $M\geq 3$, the inequality in (20) turns out to be $f(e_{S})>(2w-M)(w-1)$, which leads to a strict inequality in (21) whenever $\mathcal{C}_{3}\neq\emptyset$.
The upper bounds stated in Theorem 3 and Theorem 4 remain valid when $M=1$. So, both they can be viewed as generalizations of Theorem 2.
It is worth mentioning that the proof method used for Theorem 4 is also applicable to Theorem 3 but it differs from the original proof given in [4]. For completeness, we provide the new proof of Theorem 3 in Appendix A. Moreover, our proof in Appendix A can further reveal the fact that the equality in (14) will hold only when the codewords $S$ with either $e_{S}=1$ or $e_{S}=w$ are considered. Note that all the optimal MC-CACs obtained in [3, 4] have this property.
IV Single-Channel CACs
We first introduce the $p$-ary representation of a positive integer. given a positive integer $n$, let
$\mathbb{Z}^{\times}_{n}\triangleq\{x\in\mathbb{Z}_{n}:\,\gcd(x,n)=1\}.$
$\mathbb{Z}^{\times}_{n}$ is the set of units (i.e., invertible elements) in $\mathbb{Z}^{*}_{n}$, and thus is a multiplicative group. Note that $\mathbb{Z}^{\times}_{n}=\mathbb{Z}^{*}_{n}$ when $n$ is a prime.
Let $p$ be an odd prime and $r$ a positive integer. For $c\in\mathbb{Z}_{p^{r}}$, consider the $p$-ary representation $c=c_{0}+c_{1}p+\cdots+c_{r-1}p^{r-1}$. For $t=0,1,\ldots,r-1$, let $\mathsf{L}_{t}$ be the collection of $c\in\mathbb{Z}^{*}_{p^{r}}$ whose nonzero least significant digit in its $p$-ary representation is $p^{t}$. Obviously, $|\mathsf{L}_{t}|=(p-1)p^{r-t-1}$, and $\mathsf{L_{0}},\mathsf{L}_{1},\ldots,\mathsf{L}_{r-1}$ form a partition of $\mathbb{Z}^{*}_{p^{r}}$, i.e., $\mathbb{Z}_{p^{r}}^{*}=\mathsf{L}_{0}\uplus\mathsf{L}_{1}\uplus\cdots\uplus\mathsf{L_{r-1}}$. Integers in $\mathsf{L}_{t}$ are called in the $t$-th layer.
For a non-empty $A\subseteq\mathbb{Z}^{*}_{p}$, we arise it to a subset in $\mathbb{Z}^{*}_{p^{r}}$, for any positive integer $r$, by defining
$\mathcal{S}_{r}(A)\triangleq A_{0}\uplus A_{1}\uplus\cdots\uplus A_{r-1},$ (22)
where
$A_{t}=\{c\in\mathsf{L}_{t}:\,c_{t}\in A\}.$
$\mathcal{S}_{r}(A)$ is the collection of elements in $\mathbb{Z}^{*}_{p^{r}}$ whose nonzero least significant digit values in their $p$-ary representation are in $A$. Obviously, $|A_{t}|=|A|p^{r-1-t}$ for each $t$, and thus
$|\mathcal{S}_{r}(A)|=|A|\left(1+p+\cdots+p^{r-1}\right)=|A|\frac{p^{r}-1}{p-1}.$ (23)
The following is a useful property of the $p$-ary representation of $c\in\mathbb{Z}_{p^{r}}^{*}$, which can also be found in [5, Proposition 3].
Proposition 3 ([5]).
Let $p$ be an odd prime and $r$ be a positive integer. For $j\in\mathsf{L}_{0}$ and $c\in\mathsf{L}_{t}$, $0\leq t\leq r-1$, one has $jc\in\mathsf{L}_{t}$ and
$(jc)_{t}=j_{0}\cdot c_{t}\ (\bmod\,p).$ (24)
Note that $\mathsf{L}_{0}=\mathbb{Z}^{\times}_{p^{r}}$, which is the set of units in $\mathbb{Z}^{*}_{p^{r}}$.
IV-A Codeword Length $L=(w-1)p_{1}^{r_{1}}p_{2}^{r_{2}}\cdots p_{n}^{r_{n}}$
Given a prime $p$, an element $a\in\mathbb{Z}^{*}_{p}$ is called a quadratic residue if there exists an integer $x\in\mathbb{Z}^{*}_{p}$ such that $a=x^{2}$; otherwise, $a$ is called a quadratic non-residue. For $a\in\mathbb{Z}^{*}_{p}$, denote by $\left(\frac{a}{p}\right)=1$ if $a$ is a quadratic residue modulo $p$ and 1-1 otherwise. The notation $\left(\frac{a}{p}\right)$ is called the Legendre symbol. It can be shown (e.g., [8]) that the Legendre symbol is multiplicative:
$\left(\frac{ab}{p}\right)=\left(\frac{a}{p}\right)\left(\frac{b}{p}\right).$ (25)
This subsection is devoted to generalizing the following result, which is due to [5, Theorem 6].
Theorem 5 ([5]).
Let $w$ be a positive integer and $p$ be a prime such that $p\geq w$. If
$\left(\frac{-1}{p}\right)=-1$
and
$\left(\frac{j}{p}\right)\left(\frac{j-w+1}{p}\right)=-1,\ \forall j=1,2,\ldots,w-2,$
then for any integer $r\geq 1$,
$K((w-1)p^{r},w)=\frac{p^{r}-1}{2}.$
We first introduce the following construction.
Theorem 6.
Let $w$ be a positive integer, and let $p_{1}<p_{2}<\cdots<p_{n}$ be primes satisfying $p_{1}\geq w$. If, for each $1\leq i\leq n$, it holds that
$\left(\frac{-1}{p_{i}}\right)=-1$ (26)
and
$\left(\frac{j}{p_{i}}\right)\left(\frac{j-w+1}{p_{i}}\right)=-1,\ \forall j=1,2,\ldots,w-2,$ (27)
then for any positive integers $r_{1},\ldots,r_{n}$, there exists a code $\mathcal{C}\in\mathrm{CAC^{e}}(L,w)$ with
$|\mathcal{C}|=\frac{p_{1}^{r_{1}}\cdots p_{n}^{r_{n}}-1}{2},$
where $L=(w-1)p_{1}^{r_{1}}\cdots p_{n}^{r_{n}}$.
Proof.
For convenience, let $L^{\prime}=p_{1}^{r_{1}}\cdots p_{n}^{r_{n}}$. Then, $L=(w-1)L^{\prime}$. Since $w-1,p_{1},p_{2},\ldots,p_{n}$ are pairwise coprime, one has
$\mathbb{Z}_{L}\cong\mathbb{Z}_{w-1}\times\mathbb{Z}_{L^{\prime}}\text{ and }\mathbb{Z}_{L^{\prime}}\cong\mathbb{Z}_{p_{1}^{r_{1}}}\times\cdots\times\mathbb{Z}_{p_{n}^{r_{n}}}.$
In what follows, we use ordered pairs $(z,x)\in\mathbb{Z}_{w-1}\times\mathbb{Z}_{L^{\prime}}$ to represent elements in $\mathbb{Z}_{L}$, where $x\in\mathbb{Z}_{L^{\prime}}$ are further represented by $n$-tuples $(x_{1},\ldots,x_{n})\in\mathbb{Z}_{p_{1}^{r_{1}}}\times\cdots\times\mathbb{Z}_{p_{n}^{r_{n}}}$.
Let $Q_{i}$ be the set of quadratic residues modulo $p_{i}$. For $i=1,2,\ldots,n$ define
$\widehat{Q}_{i}\triangleq\{(0,\ldots,0,g,x_{i+1},\ldots,x_{n})$ $\in\mathbb{Z}_{p_{1}^{r_{1}}}\times\cdots\times\mathbb{Z}_{p_{n}^{r_{n}}}:\,$
$g\in\mathcal{S}_{r_{i}}(Q_{i}),x_{t}\in\mathbb{Z}_{p_{t}^{r_{t}}}\text{ for }i<t\leq n\}.$
Let
$\widehat{Q}\triangleq\biguplus_{i=1}^{n}\widehat{Q}_{i}$ (28)
be the disjoint union of $\widehat{Q}_{1},\widehat{Q}_{2},\ldots,\widehat{Q}_{n}$. As $|Q_{i}|=(p_{i}-1)/2$, by (23), we have
$|\widehat{Q}|=\sum_{i=1}^{n}\frac{p_{i}^{r_{i}}-1}{2}\prod_{j=i+1}^{n}p_{j}^{r_{j}}=\frac{p_{1}^{r_{1}}\cdots p_{n}^{r_{n}}-1}{2}.$
For $a\in\widehat{Q}$, define
$S_{a}=\{j(1,a)\in\mathbb{Z}_{w-1}\times\mathbb{Z}_{L^{\prime}}:\,j=0,1,2\ldots,w-1\}$
be the $w$-subset generated by $a$ whose set of nonzero differences is of the form
$d^{*}(S_{a})$ $=\{\pm j(1,a)\in\mathbb{Z}_{w-1}\times\mathbb{Z}_{L^{\prime}}:\,j=1,2,\ldots,w-1\}$ (29)
$=\{\pm j(1,a)\in\mathbb{Z}_{w-1}\times\mathbb{Z}_{L^{\prime}}:\,j=1,2,\ldots,w-2\}\cup\{(0,\pm(w-1)a)\}.$ (30)
Assume $a=(0,\ldots,0,g,x_{i+1},\ldots,x_{n})\in\mathbb{Z}_{p_{1}}^{r_{1}}\times\cdots\times\mathbb{Z}_{p_{n}}^{r_{n}}$, i.e., $a\in\widehat{Q}_{i}$, then the elements in $d^{*}(S_{a})$ are of the form
$\pm j(1,a)=\pm j(1,(0,\ldots,0,g,x_{i+1},\ldots,x_{n}))$
or
$(0,\pm(w-1)a)=(0,(0,\ldots,0,\pm(w-1)g,\pm(w-1)x_{i+1},\ldots,\pm(w-1)x_{n})).$
In what follows, we shall prove that $\widehat{Q}$ can generate the desired code $\mathcal{C}$, that is, $d^{*}(S_{a})$, $a\in\widehat{Q}$ are mutually disjoint. Assume $a\in\widehat{Q}_{i}$ and $b\in\widehat{Q}_{i^{\prime}}$, it is obvious that $d^{*}(S_{a})\cap d^{*}(S_{b})=\emptyset$ whenever $i\neq i^{\prime}$. So, it suffices to consider $a$ and $b$ are in the same set $\widehat{Q}_{i}$, for some $i$. Suppose to the contrary that $d^{*}(S_{a})\cap d^{*}(S_{b})\neq\emptyset$, where $a=(0,\ldots,0,g,x_{i+1},\ldots,x_{n})$ and $b=(0,\ldots,0,h,y_{i+1},\ldots,y_{n})$ for some $g,h\in\mathcal{S}_{r_{i}}(Q_{i})$. Without loss of generality generality, assume $j(1,(0,\ldots,0,g,x_{i+1},\ldots,x_{n}))$ is one of the common elements in $d^{*}(S_{a})\cap d^{*}(S_{b})$, where $1\leq j\leq w-1$. Then we have
$j(1,(0,\ldots,0,g,x_{i+1},\ldots,x_{n}))=\pm k(1,(0,\ldots,0,h,y_{i+1},\ldots,y_{n}))$ (31)
for some $k\in\{1,2,\ldots,w-1\}$. There are two cases.
Case 1: j=w1j=w-1. From (31) we have $(w-1)g=\pm(w-1)h$, which implies g=hg=-h (mod $p_{i}^{r_{i}}$) due to $g\neq h$. By considering the $p_{i}$-ary representation, both $g$ and h-h are in the same layer, say $\mathsf{L}_{t}$ for some $t$, and thus $g_{t}=(-h)_{t}$ (mod $p_{i}$). Note $(-h)_{t}=(-1)\cdot h_{t}$ (mod $p_{i}$) by Proposition 3. By (25), we have
$\left(\frac{g_{t}}{p_{i}}\right)=\left(\frac{(-h)_{t}}{p_{i}}\right)=\left(\frac{-1}{p_{i}}\right)\left(\frac{h_{t}}{p_{i}}\right).$
However, $g_{t}$ and $h_{t}$ are quadratic residues modulo $p_{i}$ by the assumption that $g,h\in\mathcal{S}_{r_{i}}(Q_{i})$ and the fact that $g,h\in\mathsf{L}_{t}$. It further implies that $\left(\frac{-1}{p_{i}}\right)=1$, which contradicts the condition in (26).
Case 2: $1\leq j\leq w-2$. From (31) we have $j(1,g)=\pm k(1,h)$ in $\mathbb{Z}_{w-1}\times\mathbb{Z}_{p_{i}^{r_{i}}}$, where $1\leq k\leq w-2$. If $j(1,g)=k(1,h)$, it follows that $g=h$, where a contradiction occurs. If j(1,g)=k(1,h)j(1,g)=-k(1,h), we have j+k=w1j+k=w-1, and then get jg=(jw+1)hjg=(j-w+1)h (mod $p_{i}^{r_{i}}$). By considering the $p_{i}$-ary representation, both $jg$ and (jw+1)h(j-w+1)h are in the same layer, say $\mathsf{L}_{t}$ for some $t$. Since $j\leq w-2<p_{i}$, we have $j=j_{0}$ and $(j-w+1)=(j-w+1)_{0}$ (mod $p_{i}$), namely, both $j$ and jw+1j-w+1 are in $\mathsf{L}_{0}$. It follows from Proposition 3 that $g,h\in\mathsf{L}_{t}$. More precisely,
$g=j^{-1}\cdot jg\in\mathsf{L}_{t}\text{ and }h=(j-w+1)^{-1}\cdot(j-w+1)h\in\mathsf{L}_{t}.$
Therefore, $g_{t},h_{t}\in Q_{i}$ by assumption. By Eq. (25) and Proposition 3,
$(jg)_{t}=((j-w+1)h)_{t}\ (\bmod\,p_{i})$
[…]
$\Rightarrow\$$\left(\frac{j}{p_{i}}\right)\left(\frac{g_{t}}{p_{i}}\right)=\left(\frac{j-w+1}{p_{i}}\right)\left(\frac{h_{t}}{p_{i}}\right)$
$\Rightarrow\$$\left(\frac{j}{p_{i}}\right)=\left(\frac{j-w+1}{p_{i}}\right),$
which contradicts the condition in (27), and the proof is completed. ∎
Remark 2.
Let $\mathcal{C}$ be the CAC constructed in Theorem 6. By (30), we have $|d^{*}(S)|=2(w-1)$ for all $S\in\mathcal{C}$. So, it costs $(w-1)(L^{\prime}-1)$ distinct differences in total, where $L^{\prime}=p_{1}^{r_{1}}\cdots p_{n}^{r_{n}}$. Moreover, since $p_{1}\geq w$, it also reveals from (30) that $(z,0)\notin d^{*}(S)$ for all $z\in\mathbb{Z}_{w-1}$ and $S\in\mathcal{C}$. This concludes that
$\bigcup_{S\in\mathcal{C}}d^{*}(S)=\mathbb{Z}_{w-1}\times\mathbb{Z}^{*}_{L^{\prime}}.$
Theorem 7.
Let $w$ be a positive integer, and let $p_{1}<p_{2}<\cdots<p_{n}$ be primes satisfying $p_{1}\geq w$. If the two conditions in (26) and (27) hold, then for any positive integers $r_{1},\ldots,r_{n}$, we have
$K\left((w-1)p_{1}^{r_{1}}\cdots p_{n}^{r_{n}},w\right)=\frac{p_{1}^{r_{1}}\cdots p_{n}^{r_{n}}-1}{2}$
provided that
$\sum_{i=1}^{t}(2w-1-p_{i})\leq w-1,$ (32)
where $p_{t}<2w-1\leq p_{t+1}$ for some $0\leq t\leq n$. Note that $p_{0}\triangleq 1$ and $p_{n+1}\triangleq\infty$ by convention. Note also that (32) always holds for $t=0,1$.
Proof.
Let $L^{\prime}=p_{1}^{r_{1}}\cdots p_{n}^{r_{n}}$. By Theorem 6, it suffices to show that for any code $\mathcal{C}\in\mathrm{CAC}((w-1)L^{\prime},w)$, one has $|\mathcal{C}|\leq\frac{L^{\prime}-1}{2}$.
Let $\mathcal{E}\subseteq\mathcal{C}$ be the collection of all exceptional codewords in $\mathcal{C}$. For notational convenience, denote by $H_{S}=\mathsf{H}(d(S))$ for $S\in\mathcal{E}$. Pick any $S\in\mathcal{E}$. Since $H_{S}$ is a subgroup of $\mathbb{Z}_{(w-1)L^{\prime}}$, the size $|H_{S}|$ must divide $(w-1)L^{\prime}$. By Lemma 1(iii), $|H_{S}|$ does not divide (w1)(w-1). As $|H_{S}|\geq 2$ by Lemma 1(i), then $|H_{S}|$ is a multiple of $p_{i}$ for some $i$. We consider two cases.
Case 1: $p_{1}>2w-2$. In this case we have $|H_{S}|\leq 2w-2<p_{1}$ due to Lemma 1(i). This is a contradiction to $|H_{S}|$ a multiple of some prime factor $p_{i}$. In other words, there is no exceptional codeword in this case. Therefore, $|d^{*}(S)|\leq 2w-2$ for all $S\in\mathcal{C}$. By the disjoint-difference-set property,
$(w-1)L^{\prime}-1=|\mathbb{Z}^{*}_{(w-1)L^{\prime}}|\geq\bigcup_{S\in\mathcal{C}}|d^{*}(S)|\geq(2w-2)|\mathcal{C}|,$
which implies that
$|\mathcal{C}|\leq\left\lfloor\frac{L^{\prime}-1}{2}+\frac{w-2}{2w-2}\right\rfloor=\frac{L^{\prime}-1}{2}.$
Case 2: $p_{1}\leq 2w-2$. Since $0\in d(S)$ for $S\in\mathcal{E}$, if follows from Proposition 1(ii) that $H_{S}\subseteq d(S)$. Then, $H_{S}\cap H_{S^{\prime}}=\{0\}$ for any two distinct $S,S^{\prime}\in\mathcal{E}$. Moreover, since $H_{S}$ is uniquely determined by its order due to Proposition 2, we have $\gcd(|H_{S}|,|H_{S^{\prime}}|)=1$. Now, assume $p_{t}<2w-1\leq p_{t+1}$ for some $1\leq t\leq n$. Since $|H_{S}|\leq 2w-2<p_{t+1}$, for $S\in\mathcal{E}$, the prime factors of $|H_{S}|$ are in $\{p_{1},\ldots,p_{t}\}$. This concludes that $|\mathcal{E}|\leq t$ and $|H_{S}|$, $S\in\mathcal{E}$, are pairwise coprime. Therefore,
$\sum_{S\in\mathcal{E}}(2w-1-|H_{S}|)\leq\sum_{i=1}^{t}(2w-1-p_{i}).$ (33)
On the other hand, by the disjoint-difference-set property,
$(w-1)L^{\prime}-1$ $=|\mathbb{Z}^{*}_{(w-1)L^{\prime}}|\geq\bigcup_{S\in\mathcal{C}}|d^{*}(S)|$
$\geq(2w-2)(|\mathcal{C}|-|\mathcal{E}|)+\sum_{S\in\mathcal{E}}|d^{*}(S)|.$
Since $|H_{S}|\leq|d(S)|=|d^{*}(S)|+1$ due to Proposition 1(ii), it follows that
$(w-1)L^{\prime}-1\geq(2w-2)|\mathcal{C}|-\sum_{S\in\mathcal{E}}(2w-1-|H_{S}|).$ (34)
Combining (32) – (34) yields
$|\mathcal{C}|\leq\left\lfloor\frac{(w-1)L^{\prime}+w-2}{2w-2}\right\rfloor=\frac{L^{\prime}-1}{2}+\left\lfloor\frac{2w-3}{2w-2}\right\rfloor=\frac{L^{\prime}-1}{2},$
as desired. ∎
The primes that satisfy the two conditions (26) and (27) are characterized in [5] with the help of Gauss’s lemma and laws of quadratic reciprocity, which lead to a series of optimal CACs of length $(w-1)p^{r}$ in [5, Corollary 2]. As an application of Theorem 7, the following corollary generalizes these optimal CACs.
Corollary 1.
Let $L^{\prime}=\prod_{i=1}^{n}p_{i}^{r_{i}}$, where $p_{i}$ are distinct primes and $r_{i}$ are any positive integers. One has
  1. 1.
    $K(3L^{\prime},4)=(L^{\prime}-1)/2$ if $p_{i}\equiv\pm 1\ (\bmod{8})$ for all $i$,
  2. 2.
    $K(4L^{\prime},5)=(L^{\prime}-1)/2$ if $p_{i}\equiv-1\ (\bmod\ 12)$ for all $i$,
  3. 3.
    $K(5L^{\prime},6)=(L^{\prime}-1)/2$ if $p_{i}\equiv-1,-5\ (\bmod\ 24)$ for all $i$,
  4. 4.
    $K(6L^{\prime},7)=(L^{\prime}-1)/2$ if $p_{i}\equiv-1,-9\ (\bmod\ 40)$ for all $i$,
  5. 5.
    $K(7L^{\prime},8)=(L^{\prime}-1)/2$ if $p_{i}\equiv-1,-49\ (\bmod\ 120)$ for all $i$,
  6. 6.
    $K(8L^{\prime},9)=(L^{\prime}-1)/2$ if $p_{i}\equiv-1,59,-109,-121,131,-169\ (\bmod\ 420)$ for all $i$,
  7. 7.
    $K(9L^{\prime},10)=(L^{\prime}-1)/2$ if $p_{i}\equiv-1,-9,31,-81,111,-121\ (\bmod\ 280)$ for all $i$, and
  8. 8.
    $K(10L^{\prime},11)=(L^{\prime}-1)/2$ if $p_{i}\equiv-1,-5,-25,43,47,67\ (\bmod\ 168)$ for all $i$.
Note that we can derive a sufficient condition of primes $p_{i}$ so that $K((w-1)L^{\prime},w)=(L^{\prime}-1)/2$, for any arbitrary $w$. Please refer to [5] for more details.
IV-B Codeword Length $L=p_{1}^{r_{1}}p_{2}^{r_{2}}\cdots p_{n}^{r_{n}}$
This subsection is devoted to generalize the following result, which is due to [5, Theorem 10].
Theorem 8 ([5]).
Let $w$ be a positive integer, and let $p$ be a prime such that p1p-1 is divided by 2w22w-2. If there is a code in $\mathrm{CAC^{e}}(p,w)$ with (p1)/(2w2)(p-1)/(2w-2) codewords, then for any integer $r\geq 1$,
$K\left(p^{r},w\right)=\frac{p^{r}-1}{2w-2}.$
We first introduce the following construction.
Theorem 9.
Let $w$ be a positive integer, and let $p_{1}<p_{2}<\cdots<p_{n}$ be primes satisfying $p_{1}\geq 2w-1$. If there is a code in $\mathrm{CAC^{e}}(p_{i},w)$ with $m_{i}$ codewords for each $i$, then for any positive integers $r_{1},\ldots,r_{n}$, there exists a code $\mathcal{C}\in\mathrm{CAC^{e}}(L,w)$ with
$|\mathcal{C}|=\sum_{i=1}^{n}\frac{m_{i}(p_{i}^{r_{i}}-1)}{p_{i}-1}\prod_{j=i+1}^{n}p_{j}^{r_{j}},$ (35)
where $L=p_{1}^{r_{1}}\cdots p_{n}^{r_{n}}$.
Proof.
Since $p_{1},p_{2},\ldots,p_{n}$ are distinct primes,
$\mathbb{Z}_{L}\cong\mathbb{Z}_{p_{1}^{r_{1}}}\times\mathbb{Z}_{p_{2}^{r_{2}}}\times\cdots\times\mathbb{Z}_{p_{n}^{r_{n}}}.$
In what follows, we use $n$-tuples $(x_{1},\ldots,x_{n})\in\mathbb{Z}_{p_{1}^{r_{1}}}\times\cdots\times\mathbb{Z}_{p_{n}^{r_{n}}}$ to represent elements in $\mathbb{Z}_{L}$.
For $i=1,2,\ldots,n$, let $\Gamma_{i}$ be a set of $m_{i}$ generators of the given code in $\mathrm{CAC^{e}}(p_{i},w)$. Then, define
[…]
$g\in\mathcal{S}_{r_{i}}(\Gamma_{i}),x_{i}\in\mathbb{Z}_{p_{i}}^{r_{i}}\text{ for }i<j\leq n\}.$
It is easy to see from (23) that for $i=1,2,\ldots,n$,
$|\widehat{\Gamma}_{i}|=\frac{m_{i}(p_{i}^{r_{i}}-1)}{p_{i}-1}\prod_{j=i+1}^{n}p_{j}^{r_{j}}.$
Note that the sets $\widehat{\Gamma}_{1},\widehat{\Gamma}_{2},\ldots,\widehat{\Gamma}_{n}$ are mutually disjoint. Let
$\widehat{\Gamma}=\biguplus_{i=1}^{n}\widehat{\Gamma}_{i}$ (36)
be the disjoint union of these sets. Obviously,
$|\widehat{\Gamma}|=\sum_{i=1}^{n}\frac{m_{i}(p_{i}^{r_{i}}-1)}{p_{i}-1}\prod_{j=i+1}^{n}p_{j}^{r_{j}}.$
Consider the set $\widehat{\Gamma}$. For $a\in\widehat{\Gamma}$, if $a=(0,\ldots,0,g,x_{i+1},\ldots,x_{n})$ for some $g\in\mathcal{S}_{r_{i}}(\Gamma_{i})$, i.e., $a\in\widehat{\Gamma}_{i}$, then define
$S_{a}=\{ja\in\mathbb{Z}_{p_{1}^{r_{1}}}\times\cdots\times\mathbb{Z}_{p_{n}^{r_{n}}}:\,j=0,1,2,\ldots,w-1\}$ (37)
be the $w$-subset generated by $a$ whose set of nonzero differences is of the form
$d^{*}(S_{a})=\{ja\in\mathbb{Z}_{p_{1}^{r_{1}}}\times\cdots\times\mathbb{Z}_{p_{n}^{r_{n}}}:\,j=\pm 1,\pm 2,\ldots,\pm(w-1)\}.$ (38)
In what follows, we shall prove that $\widehat{\Gamma}$ is the set of generators of the desired code $\mathcal{C}$, that is, $d^{*}(S_{a})\cap d^{*}(S_{b})=\emptyset$ for any distinct $a,b\in\widehat{\Gamma}$. It suffices to consider that $a$ and $b$ are in the same subset $\widehat{\Gamma}_{i}$, for some $i$, because the assertion is definitely true in the other case.
Suppose to the contrary that $d^{*}(S_{a})\cap d^{*}(S_{b})\neq\emptyset$, where $a=(0,\ldots,0,g,x_{i+1},\ldots,x_{n})$ and $b=(0,\ldots,0,h,y_{i+1},\ldots,y_{n})$ for some $g,h\in\mathcal{S}_{r_{i}}(\Gamma_{i})$. Assume $j(0,\ldots,0,g,x_{i+1},\ldots,x_{n})=k(0,\ldots,0,h,y_{i+1},\ldots,y_{n})$ for some $j,k\in\{\pm 1,\pm 2,\ldots,\pm(w_{i}-1)\}$. If $j=k$, then $g=h$ and $x_{t}=y_{t}$ for $i+1\leq t\leq n$, namely, $a=b$. So, we consider $j\neq k$ in what follows. Let us focus on the $i$-th component, namely, $jg=kh$. Since $p_{i}\geq 2w-1$, by Proposition 3, one has $jg\in\mathsf{L}_{t}$ for every $j\in\{\pm 1,\pm 2,\ldots,\pm(w-1)\}$ if $g\in\mathsf{L}_{t}$. Here we use $\mathsf{L}_{t}$ to denote the $t$-th layer in the $p_{i}$-ary representation of elements of $\mathbb{Z}_{p_{i}^{r_{i}}}$ for notatinoal convenience. It follows that $jg\neq kh$ for any $j,k\in\{\pm 1,\pm 2,\ldots,\pm(w-1)\}$ whenever $g$ and $h$ are in distinct layers. Therefore, we only need to consider the case when both $g$ and $h$ are in the same layer, say $\mathsf{L}_{t}$ for some $t$. Observe that $j,k\in\mathsf{L}_{0}$ due to $p_{i}\geq 2w-1$. By Proposition 3 again, $j\cdot g_{t}=k\cdot h_{t}$ (mod $p_{i}$). If $g_{t}\neq h_{t}$, by the assumption that $g_{t},h_{t}\in\Gamma_{i}$ be two distinct generators in the given code in $\mathrm{CAC^{e}}(p_{i},w)$, one has $j\cdot g_{t}\neq k\cdot h_{t}$. If $g_{t}=h_{t}$, it further implies that $(j-k)g_{t}=0$ (mod $p_{t}$), which is impossible because of $j,k$ beging distinct in $\{\pm 1,\pm 2,\cdots,\pm(w-1)\}$ and $p_{i}\geq 2w-1$. This completes the proof. ∎
Theorem 10.
Let $w$ be a positive integer, and let $p_{1}<\cdots<p_{n}$ be primes such that $p_{i}-1$ is divided by 2w22w-2 for each $i$. If there is a code in $\mathrm{CAC^{e}}(p_{i},w)$ with $(p_{i}-1)/(2w-2)$ codewords for each $i$, then for any positive integers $r_{1},\ldots,r_{n}$,
$K(p_{1}^{r_{1}}\cdots p_{t}^{r_{n}},w)=\frac{p_{1}^{r_{1}}\cdots p_{t}^{r_{t}}-1}{2(w-1)}.$ (39)
Proof.
The assumption that $p_{i}-1$ is divided by 2w22w-2 guarantees $p_{i}\geq 2w-1$, for each $i$. By plugging $m_{i}=(p_{i}-1)/(2w-2)$ into (35), there exists a code in $\mathrm{CAC^{e}}(p_{1}^{r_{1}}\cdots p_{n}^{r_{n}},w)$ with $(p_{1}^{r_{1}}\cdots p_{n}^{r_{n}}-1)/(2w-2)$ codewords. Since the least prime factor $p_{1}$ is larger than or equal to 2w12w-1, the result follows by Theorem 2. ∎
Remark 3.
Let $\mathcal{C}$ be an optimal CAC considered in Theorem 10. Since $p_{1}\geq 2w-1$, it follows from (38) that $|d^{*}(S)|=2(w-1)$ for $S\in\mathcal{C}$. As $|\mathcal{C}|=(L-1)/2$, where $L=p_{1}^{r_{1}}\cdots p_{n}^{r_{n}}$, we have
$\bigcup_{S\in\mathcal{C}}d^{*}(S)=\mathbb{Z}^{*}_{L}.$
So, the optimal CACs obtained in Theorem 9 is tight [9].
V Multichannel CACs with Two Channels
Theorem 11.
Let $w$ be a positive integer, and let $p_{1}<p_{2}<\cdots<p_{n}$ be primes satisfying $p_{1}\geq w$. For each $1\leq i\leq n$, if there is a code in $\mathrm{CAC^{e}}(p_{i},w)$ with $m_{i}$ codewords and the two conditions in (26) and (27) hold, then for any positive integers $r_{1},\ldots,r_{n}$, there exists a code $\mathcal{C}\in\text{MC-CAC}(2,L,w)$ with
$|\mathcal{C}|=\prod_{i=1}^{n}p_{i}^{r_{i}}+\sum_{i=1}^{n}\frac{2m_{i}(p_{i}^{r_{i}}-1)}{p_{i}-1}\prod_{j=i+1}^{n}p_{j}^{r_{j}}$ (40)
codewords, where $L=(w-1)p_{1}^{r_{1}}p_{2}^{r_{2}}\cdots p_{n}^{r_{n}}$.
Proof.
Let $L^{\prime}=p_{1}^{r_{1}}p_{2}^{r_{2}}\cdots p_{n}^{r_{n}}$, then $L=(w-1)L^{\prime}$. In the following construction, the desired code $\mathcal{C}$ will consist of three classes of codewords $\mathcal{C}_{1}$, $\mathcal{C}_{2}$ and $\mathcal{C}_{3}$ with sizes
$|\mathcal{C}_{1}|=L^{\prime}-1,\ |\mathcal{C}_{2}|=\sum_{i=1}^{n}\frac{2m_{i}(p_{i}^{r_{i}}-1)}{p_{i}-1}\prod_{j=i+1}^{n}p_{j}^{r_{j}},\ \text{and }|\mathcal{C}_{3}|=1,$
where the codewords in $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are from the constructions in Theorem 6 and Theorem 9, respectively.
Since $\gcd(w-1,L^{\prime})=1$, we have $\mathbb{Z}\cong\mathbb{Z}_{w-1}\times\mathbb{Z}_{L^{\prime}}$. Elements in $\mathcal{I}_{2}\times\mathbb{Z}_{L}$ can be represented by $(m,(z,x))$, where $m\in\mathcal{I}_{2}=\{1,2\}$ and $z\in\mathbb{Z}_{w-1}$, $x\in\mathbb{Z}_{L^{\prime}}$.
Since the two conditions (26) – (27) hold, Theorem 6 ensures the existence of a code in $\mathrm{CAC}(L,w)$ with $(L^{\prime}-1)/2$ codewords. We adopt the notation in the proof of Theorem 6. Consider the set $\widehat{Q}$ defined in (28). Note that $\widehat{Q}\subset\mathbb{Z}_{L^{\prime}}$ with size $|\widehat{Q}|=(L^{\prime}-1)/2$. For $a\in\widehat{Q}$ define the following two codewords in $\mathcal{I}_{2}\times\mathbb{Z}_{L}$
$S_{a}^{1}$ $=\{(1,(1,-a))\}\cup\{(2,j(1,a)):\,j=0,1,2,\ldots,w-2\},\text{ and }$
$S_{a}^{2}$ $=\{(1,j(1,a)):\,j=0,1,2,\ldots,w-2\}\cup\{(2,(1,-a))\}.$
Let $\mathcal{C}_{1}=\{S_{a}^{1},S_{a}^{2}:\,a\in\widehat{Q}\}$ be the collection of these codewords derived from all elements in $\widehat{Q}$. Let us focus on the arrays of differences $D_{S_{a}^{1}}$ and $D_{S_{a}^{2}}$. It is easy to see that
$D_{S_{a}^{1}}(1,1)=D_{S_{a}^{2}}(2,2)$ $=\emptyset,$ (41)
$D_{S_{a}^{1}}(2,2)=D_{S_{a}^{2}}(1,1)$ $=\{\pm j(1,a)\in\mathbb{Z}_{w-1}\times\mathbb{Z}_{L^{\prime}}:\,j=1,2,\ldots,w-2\},$ (42)
$D_{S_{a}^{1}}(1,2)=D_{S_{a}^{2}}(2,1)$ $=\{-j(1,a)\in\mathbb{Z}_{w-1}\times\mathbb{Z}_{L^{\prime}}:\,j=1,2,\ldots,w-1\},$ (43)
$D_{S_{a}^{1}}(2,1)=D_{S_{a}^{2}}(1,2)$ $=\{j(1,a)\in\mathbb{Z}_{w-1}\times\mathbb{Z}_{L^{\prime}}:\,j=1,2,\ldots,w-1\}.$ (44)
Recall that in the proof of Theorem 6, it was shown that $d^{*}(S_{a})\cap d^{*}(S_{b})=\emptyset$ for any two distinct $a,b\in\widehat{Q}$, where $d^{*}(S_{a})$ is given in (29) as
$d^{*}(S_{a})=\{\pm j(1,a)\in\mathbb{Z}_{w-1}\times\mathbb{Z}_{L^{\prime}}:\,j=1,2,\ldots,w-1\}.$
Therefore, it follows from (41) – (44) that $D_{S_{a}^{1}},D_{S_{a}^{2}}$ are mutually disjoint over all $a\in\widehat{Q}$. Moreover, by the property given in Remark 2, we further derive that
[…]
Since there is a code in $\mathrm{CAC^{e}}(p_{i},w)$ with $m_{i}$ codewords for each $i$, it follows from Theorem 9 that there exists a code in $\mathrm{CAC}(L^{\prime},w)$ with $\sum_{i=1}^{n}\left(\frac{m_{i}(p_{i}^{r_{i}}-1)}{p_{i}-1}\prod_{j=i+1}^{n}p_{j}^{r_{j}}\right)$ codewords, where the generators are collected in the set $\widehat{\Gamma}$ given in (36). We adopt the notation in the proof of Theorem 9. For $a\in\widehat{\Gamma}$ define the following two codewords in $\mathcal{I}_{2}\times\mathbb{Z}_{L^{\prime}}$
$T_{a}^{1}$ $=\{(1,(0,ja)):\,j=0,1,2,\ldots,w-1\},\text{ and }$
$T_{a}^{2}$ $=\{(2,(0,ja)):\,j=0,1,2,\ldots,w-1\}.$
Let $\mathcal{C}_{2}=\{T_{a}^{1},T_{a}^{2}:\,a\in\widehat{\Gamma}\}$ be the collection of these codewords. Consider the arrays of differences $D_{T_{a}^{1}}$ and $D_{T_{a}^{2}}$. It is obvious that
$D_{T_{a}^{1}}(1,2)=D_{T_{a}^{1}}(2,1)=D_{T_{a}^{1}}(2,2)=D_{T_{a}^{2}}(1,2)=D_{T_{a}^{2}}(2,1)=D_{T_{a}^{2}}(1,1)=\emptyset,\text{ and}$ (48)
$D_{T_{a}^{1}}(1,1)=D_{T_{a}^{2}}(2,2)=\{(0,\pm ja)\in\mathbb{Z}_{w-1}\times\mathbb{Z}_{L^{\prime}}:\,j=1,2,\ldots,w-1\}.$ (49)
In the proof of Theorem 9, it was shown that $\{\pm ja\in\mathbb{Z}_{L^{\prime}}:\,j=1,2,\ldots,w-1\}$, $a\in\widehat{\Gamma}$, are mutually disjoint. Then, it follows from (48) – (49) that $D_{T_{a}^{1}},D_{T_{a}^{2}}$ are mutually disjoint over all $a\in\widehat{\Gamma}$. Moreover, we have
[…]
Finally, the third class $\mathcal{C}_{3}$ only contains one codeword given by
$U=\{(1,(j,0)):\,j=0,1,\ldots,w-2\}\cup\{(2,(0,0))\}.$
Observe that
$D_{U}=\begin{tabular}[]{|c|c|}\hline\cr$\mathbb{Z}^{*}_{w-1}\times\{0\}$&$\mathbb{Z}_{w-1}\times\{0\}$\\ \hline\cr$\mathbb{Z}_{w-1}\times\{0\}$$\emptyset$\\ \hline\cr\end{tabular}$. (55)
By (47), (52), (55) and aforementioned arguments, $\mathcal{C}_{1}\cup\mathcal{C}_{2}\cup\mathcal{C}_{3}$ forms the desired MC-CAC. ∎
We have the following upper bound on the value $K(2,L,w)$ when the length $L$ is a product of larger prime factors.
Theorem 12.
Let $L^{\prime}$ and $w$ be positive integers. If the prime factors of $L^{\prime}$ are all larger than or equal to 2w12w-1, then
$K(2,(w-1)L^{\prime},w)\leq L^{\prime}+\left\lfloor\frac{L^{\prime}+w-3}{w-1}\right\rfloor.$
Proof.
Let $L=(w-1)L^{\prime}$, and let $\mathcal{C}$ be any code in $\text{MC-CAC}(2,L,w)$. We first partition codewords in $\mathcal{C}$ into two classes, $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$, by
$\mathcal{C}_{1}$ $=\{S\in\mathcal{C}:\,e_{S}=1\},$
$\mathcal{C}_{2}$ $=\{S\in\mathcal{C}:\,e_{S}=2\}.$
Consider the codewords in $\mathcal{C}_{1}$. Suppose $S\in\mathcal{C}_{1}$ is exceptional. By Lemma 1(i), we have $|\mathsf{H}(S-S)|\leq 2w-2$, which is strictly less than any factor of $L^{\prime}$. Note that $|\mathsf{H}(S-S)|$ divides $L=(w-1)L^{\prime}$ as $\mathsf{H}(S-S)$ is a subgroup of $\mathbb{Z}_{L}$ due to Proposition 1(i). It must be the case that $|\mathsf{H}(S-S)|$ divides w1w-1. This is a contradiction to Lemma 1(iii). Therefore, there is no exceptional codeword in $\mathcal{C}_{1}$. It follows that
$\sum_{S\in\mathcal{C}_{1}}|D_{S}(1,1)|+|D_{S}(2,2)|\geq(2w-2)|\mathcal{C}_{1}|.$ (56)
Now, consider the codewords in $\mathcal{C}_{2}$. If $\{S_{1},S_{2}\}$ is an exceptional pair, by Lemma 2(i), we have $|\mathsf{H}(S_{1}-S_{2})|\leq|S_{1}|+|S_{2}|-2=w-2$, which is strictly less than any factor of $L^{\prime}$. As $\mathsf{H}(S_{1}-S_{2})$ is a subgroup of $\mathbb{Z}_{L}$, it implies that $|\mathsf{H}(S_{1}-S_{2})|$ divides w1w-1, which is a contradiction to Lemma 2(ii). So, $\{S_{1},S_{2}\}$ is not an exceptional pair. Hence, for $S\in\mathcal{C}_{2}$, one has $|D_{S}(1,2)|=|S_{1}-S_{2}|\geq|S_{1}|+|S_{2}|-1=w-1$. It follows that
$(w-1)L^{\prime}\geq\sum_{S\in\mathcal{C}}|D_{S}(1,2)|=\sum_{S\in\mathcal{C}_{2}}|D_{S}(1,2)|\geq(w-1)|\mathcal{C}_{2}|.$ (57)
Following the arguments above, if $S$ is an exceptional codeword in $\mathcal{C}_{2}$, it must be the case that either $S_{1}$, $S_{2}$, or both are exceptional. To make a more accurate estimation of the number of differences that are derived from all exceptional codewords, we define, for $i=1,2$,
$\mathcal{C}_{2}^{(i)}\triangleq\{S_{i}:\,S\in\mathcal{C}_{2}\}$
and
$\mathcal{E}^{(i)}\triangleq\{S_{i}\in\mathcal{C}_{2}^{(i)}:\,S_{i}\text{ is exceptional}\}.$
Note that $|\mathcal{C}_{2}^{(1)}|=|\mathcal{C}_{2}^{(2)}|=|\mathcal{C}_{2}|$ and
$\sum_{A\in\mathcal{C}_{2}^{(1)}}|A|+\sum_{B\in\mathcal{C}_{2}^{(1)}}|B|=w|\mathcal{C}_{2}|.$ (58)
Consider an element $A\in\mathcal{E}^{(1)}$. For notational convenience, denote by $H_{A}=\mathsf{H}(A-A)$. Since, by Proposition 1(i), $H_{A}$ contains the element 0, so we can further denote by $H_{A}^{*}=H_{A}\setminus\{0\}$. By the same argument given above, $|H_{A}|$ must divide w1w-1. As $H_{A}$ is a subgroup of $\mathbb{Z}_{L}$, we have $H_{A}=-H_{A}$, which implies that $|-A+H_{A}|=|-(A+H_{A})|=|A+H_{A}|$. By plugging $A=A$ and B=AB=-A into (6), we have
$|d(A)|=|A+(-A)|$ $\geq|A+H_{A}|+|-A+H_{A}|-|H_{A}|$
$=2|A+H_{A}|-|H_{A}|\geq 2|A|-|H_{A}^{*}|-1,$
which leads to
$|d^{*}(A)|\geq 2|A|-|H^{*}_{A}|-2.$ (59)
Since $H_{A}$ is a subgroup of $\mathbb{Z}_{L}$ and $|H_{A}|$ divides w1w-1, by Proposition 2, $H_{A}$ is a subgroup of $G=\{iL^{\prime}:\,i=0,1,\ldots,w-2\}$. Moreover, observe that $H_{A}\subseteq d(A)$ due to Proposition 1(ii). Then, by the definition of an MC-CAC, $H^{*}_{A}\cap H^{*}_{A^{\prime}}=\emptyset$ for any pair of distinct $A,A^{\prime}\in\mathcal{E}^{(1)}$. This concludes that
$\sum_{A\in\mathcal{E}^{(1)}}|H^{*}_{A}|=\left|\biguplus_{A\in\mathcal{E}^{(1)}}H^{*}_{A}\right|\leq|G\setminus\{0\}|=w-2.$
By (59), we further have
$\sum_{A\in\mathcal{E}^{(1)}}|d^{*}(A)|\geq\left(\sum_{A\in\mathcal{E}^{(1)}}2|A|-2\right)-(w-2).$
Since $|d^{*}(A)|\geq 2|A|-2$ if $A$ is not exceptional, it follows that
$\sum_{A\in\mathcal{C}_{2}^{(1)}}|d^{*}(A)|\geq\left(\sum_{A\in\mathcal{C}_{2}^{(1)}}2|A|-2\right)-(w-2).$ (60)
Similarly, we have
$\sum_{B\in\mathcal{C}_{2}^{(2)}}|d^{*}(B)|\geq\left(\sum_{B\in\mathcal{C}_{2}^{(2)}}2|B|-2\right)-(w-2).$ (61)
By (56), (58), (60) and (61),
$2(w-1)L^{\prime}-2=2|\mathbb{Z}^{*}_{L}|$ $\geq\sum_{S\in\mathcal{C}}|D_{S}(1,1)|+\sum_{S\in\mathcal{C}}|D_{S}(2,2)|$
$=\sum_{S\in\mathcal{C}_{1}}\left(|D_{S}(1,1)|+|D_{S}(2,2)|\right)+\sum_{S\in\mathcal{C}_{2}}\left(|D_{S}(1,1)|+|D_{S}(2,2)|\right)$
$\geq(2w-2)|\mathcal{C}_{1}|+\sum_{A\in\mathcal{C}_{2}^{(1)}}|d^{*}(A)|+\sum_{B\in\mathcal{C}_{2}^{(1)}}|d^{*}(B)|$
$\geq(2w-2)|\mathcal{C}_{1}|+(2w-4)|\mathcal{C}_{2}|-2(w-2),$
which yields
$(w-1)L^{\prime}-1\geq(w-1)|\mathcal{C}_{1}|+(w-2)|\mathcal{C}_{2}|-(w-2).$ (62)
Combining it with (57), we obtain
$wL^{\prime}-1\geq(w-1)|\mathcal{C}|-(w-2),$
and hence the result follows. ∎
We now show that the construction described in Theorem 11 is optimal.
Theorem 13.
Let $w$ be positive integer, and let $p_{1}<\cdots<p_{n}$ be primes such that $p_{i}-1$ is divisible by 2w22w-2 for each $i$. For each $1\leq i\leq n$, if there is a code in $\mathrm{CAC^{e}}(p_{i},w)$ with $(p_{i}-1)/(2w-2)$ codewords and the two conditions in (26) and (27) hold, then for any positive integers $r_{1},\ldots,r_{n}$,
$K(2,L,w)=\prod_{i=1}^{n}p_{i}^{r_{i}}+\frac{\left(\prod_{i=1}^{n}p_{i}^{r_{i}}\right)-1}{w-1},$
where $L=(w-1)p_{1}^{r_{1}}\cdots p_{n}^{r_{n}}$.
Proof.
The assumption that $p_{1}-1$ is divisible by 2w22w-2 guarantees that $p_{1}\geq 2w-1\geq w$. By plugging $m_{i}=(p_{i}-1)/(2w-2)$ into Theorem 11, there exists a code in $\text{MC-CAC}(2,L,w)$ with $\prod_{i=1}^{n}p_{i}^{r_{i}}+\left(\left(\prod_{i=1}^{n}p_{i}^{r_{i}}\right)-1\right)/(w-1)$ codewords.
Let $\mathcal{C}$ be any code in $\text{MC-CAC}(2,L,w)$. By plugging $L^{\prime}=\prod_{i=1}^{n}p_{i}^{r_{i}}$ into Theorem 12,
$|\mathcal{C}|$ $\leq\prod_{i=1}^{n}p_{i}^{r_{i}}+\frac{\left(\prod_{i=1}^{n}p_{i}^{r_{i}}\right)-1}{w-1}+\left\lfloor\frac{w-2}{w-1}\right\rfloor$
$=\prod_{i=1}^{n}p_{i}^{r_{i}}+\frac{\left(\prod_{i=1}^{n}p_{i}^{r_{i}}\right)-1}{w-1}.$
This completes the proof. ∎
VI Multichannel CACs with $M>2$ Channels
Theorem 14.
Let $M,L,L^{\prime}$, and $w$ be positive integers such that $w>M\geq 3$, $M$ divides $w$, $L=(2\frac{w}{M}-1)L^{\prime}$, and all prime factors of $L^{\prime}$ are larger than or equal to 2w12w-1. If there is a code in $\mathrm{CAC^{e}}(L^{\prime},w)$ with $(L^{\prime}-1)/(2w-2)$ codewords, then there exists a code $\mathcal{C}\in\text{MC-CAC}(M,L,w)$ with
$|\mathcal{C}|=\frac{M(M-1)L}{(2w-M)(w-1)}+\frac{M(L-1)}{2w-2}.$
Proof.
For convenience, let $t=\frac{w}{M}$. Then $L=(2t-1)L^{\prime}$. Since every prime factor $p$ of $L^{\prime}$ satisfies $p\geq 2w-1>2t-1$, then $\gcd(2t-1,L^{\prime})=1$, thus $\mathbb{Z}_{L}\cong\mathbb{Z}_{2t-1}\times\mathbb{Z}_{L^{\prime}}$. So, for the sake of conveniences, the elements in $\mathbb{Z}_{L}$ can be represented as ordered pairs in $\mathbb{Z}_{2t-1}\times\mathbb{Z}_{L^{\prime}}$.
The objective code will be decomposed into two classes, $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$, whose constructions are given below.
Let $\mathcal{C}^{\prime}$ be the code in $\mathrm{CAC}(L^{\prime},w)$ with $|\mathcal{C}^{\prime}|=(L^{\prime}-1)/2(w-1)$ codewords. For $S\in\mathcal{C}^{\prime}$ and $m\in\mathcal{I}_{M}$, define a $w$-subset in $\mathcal{I}_{M}\times\mathbb{Z}_{L}$ by
$\widehat{S}_{m}\triangleq\{(m,(0,a))\in\mathcal{I}_{M}\times\mathbb{Z}_{L}:\,a\in S\}.$ (63)
Let
$\mathcal{C}_{1}=\bigcup_{S\in\mathcal{C}^{\prime},\,m\in\mathcal{I}_{M}}\widehat{S}_{m}.$
Obviously, $|\mathcal{C}_{1}|=M(L^{\prime}-1)/(2w-2)$. It is easy to see that
$D_{\widehat{S}_{m}}(i,j)=\begin{cases}\{0\}\times d^{*}(S)&\text{if }i=j=m,\\ \emptyset&\text{otherwise}.\end{cases}$ (64)
As $\mathcal{C}^{\prime}$ is a CAC, it holds that $D_{\widehat{S}_{m}}$, $\widehat{S}_{m}\in\mathcal{C}_{1}$, are mutually disjoint.
Secondly, let $\Omega=\{(1,g)\in\mathbb{Z}_{2t-1}\times\mathbb{Z}_{L^{\prime}}:\,g\in\mathbb{Z}_{L^{\prime}}\}\ (=\{1\}\times\mathbb{Z}_{L^{\prime}})$. For $a\in\Omega$, define a $w$-subset in $\mathcal{I}_{M}\times\mathbb{Z}_{L}$ by
$S_{a}\triangleq\bigcup_{m=1}^{M}\{(m,ka)\in\mathcal{I}_{M}\times\mathbb{Z}_{L}:\,k=(m-1)t,(m-1)t+1,\ldots,mt-1\}.$ (65)
Let
$\mathcal{C}_{2}=\bigcup_{a\in\Omega}S_{a}.$
It is not difficult to see that
$D_{S_{a}}(i,j)=\begin{cases}\{ka\in\mathbb{Z}_{L}:\,k=\pm 1,\pm 2,\ldots,\pm(t-1)\}&\text{if }i=j,\\ \{ka\in\mathbb{Z}_{L}:\,(i-j-1)t+1\leq k\leq(i-j+1)t-1\}&\text{if }i\neq j.\end{cases}$ (66)
For convenience, let
$A_{i,j}=\{x\in\mathbb{Z}:\,(i-j-1)t+1\leq x\leq(i-j+1)t-1\}.$
We will claim that $D_{S_{a}}\cap D_{S_{b}}=\emptyset$ for any two distinct $a,b\in\Omega$. Suppose to the contrary that $D_{S_{a}}\cap D_{S_{b}}\neq\emptyset$, where $a=(1,g),b=(1,h)$ for some distinct $g,h\in\mathbb{Z}_{L^{\prime}}$. There are two cases according to (65).
Case 1. If $D_{S_{a}}(i,i)\cap D_{S_{b}}(i,i)\neq\emptyset$ for some $i$, then there exist $k,k^{\prime}\in\{\pm 1,\pm 2,\ldots,\pm(t-1)\}$ such that $k(1,g)=k^{\prime}(1,h)$ in $\mathbb{Z}_{2t-1}\times\mathbb{Z}_{L^{\prime}}$. This implies that $k=k^{\prime}$ and thus k(gh)=0k(g-h)=0 (mod $L^{\prime}$). Since all prime factors of $L^{\prime}$ are larger than or equal to 2w12w-1, which is strictly larger than 2t12t-1, we have $\gcd(k,L^{\prime})=1$. This further implies that $g=h$, which is a contradiction.
Case 2. If $D_{S_{a}}(i,j)\cap D_{S_{b}}(i,j)\neq\emptyset$ for some $i\neq j$, then there exist integers $k,k^{\prime}\in A_{i,j}$ such that $k(1,g)=k^{\prime}(1,h)$ in $\mathbb{Z}_{2t-1}\times\mathbb{Z}_{L^{\prime}}$. Observe that $A_{i,j}$ is consist of consecutive 2t12t-1 integers. Then, we get $k=k^{\prime}$, and thus k(gh)=0k(g-h)=0 (mod $L^{\prime}$). Moreover, since $i,j$ are distinct elements in $\mathcal{I}_{M}$, we have $w+1\leq k\leq w+1$ and $k\neq 0$. Again, since all prime factors of $L^{\prime}$ are larger than or equal to 2w12w-1, one has $\gcd(k,L^{\prime})=1$. Hence, it follows that $g=h$, and a contradiction occurs.
Furthermore, by (64) and (66), it is obvious that $D_{S}\cap D_{S^{\prime}}=\emptyset$ whenever $S\in\mathcal{C}_{1}$ and $S^{\prime}\in\mathcal{C}_{2}$. It concludes that $\mathcal{C}=\mathcal{C}_{1}\uplus\mathcal{C}_{2}$ is a code in $\text{MC-CAC}(M,L,w)$. Finally, as $L^{\prime}=ML/(2w-M)$, the code size is
$|\mathcal{C}|$ $=|\mathcal{C}_{1}|+|\mathcal{C}_{2}|=\frac{M(L^{\prime}-1)}{2w-2}+L^{\prime}$
$=\frac{M+2w-2}{2w-2}\cdot\frac{ML}{2w-M}-\frac{M}{2w-2}$
$=\frac{M(M-1)L}{(w-1)(2w-M)}+\frac{M(L-1)}{2w-2},$
as desired. ∎
For any positive integer $n$, let $\tau(n)$, called the number-of-divisor function, denote the number of positive divisors of $n$.
Theorem 15.
Let $M,L,L^{\prime}$ and $w$ be positive integers with $w>M\geq 3$ such that $M$ divides $w$, $L=(2\frac{w}{M}-1)L^{\prime}$, and all prime factors of $L^{\prime}$ are larger than or equal to 2w12w-1. Then,
$K(M,L,w)\leq\frac{M(M-1)(L+(\tau(2\frac{w}{M}-1)-1)(L-L^{\prime}))}{(2w-M)(w-1)}+\frac{M(L-1)+2(w-M)}{2w-2}.$
Proof.
Let $\mathcal{C}$ be any code in $\text{MC-CAC}(M,L,w)$. We adopt the notation from the proof of Theorem 4 for convenience. Let $\mathcal{C}_{1}=\{S\in\mathcal{C}:\,e_{S}=1\}$, $\mathcal{C}_{2}=\{S\in\mathcal{C}:\,e_{S}=M\}$, and $\mathcal{C}_{3}=\{S\in\mathcal{C}:\,2\leq e_{S}\leq M-1\}$.
For $i\in\mathcal{I}_{M}$, let
$\mathcal{C}^{(i)}\triangleq\{S_{i}:\,S\in\mathcal{C}\}$
and
$\mathcal{E}^{(i)}\triangleq\{A\in\mathcal{C}^{(i)}:\,A\text{ is exceptional}\}$
denote the collections of all subsets and all exceptional subsets associated with the $i$-th channel, respectively. For $A\in\mathcal{E}^{(i)}$, let $H_{A}=\mathsf{H}(A-A)$ for notational convenience.
For any fixed index $i\in\mathcal{I}_{M}$, by the same arguments used to obtain (60), we derive
$\sum_{A\in\mathcal{E}^{(i)}}|d^{*}(A)|\geq\left(\sum_{A\in\mathcal{E}^{(i)}}2|A|-2\right)-\left(2\frac{w}{M}-2\right).$
Note that $|d^{*}(A)|\geq 2|A|-2$ for all $A\in\mathcal{C}^{(i)}\setminus\mathcal{E}^{(i)}$. Therefore,
$\sum_{A\in\mathcal{C}^{(i)}}|d^{*}(A)|\geq\left(\sum_{A\in\mathcal{C}^{(i)}}2|A|-2\right)-2\left(\frac{w}{M}-1\right).$ (67)
Summing (67) over all $i\in\mathcal{I}_{M}$, by the disjoint-difference-array property, yields
$M(L-1)\geq 2(w-1)|\mathcal{C}_{1}|+2(w-M)|\mathcal{C}_{2}|+2\sum_{S\in\mathcal{C}_{3}}(w-e_{S})-2M\left(\frac{w}{M}-1\right).$ (68)
Now, we consider $D_{S}(i,j)$ ($S_{i}-S_{j}$) for distinct indices $i,j\in\mathcal{I}_{M}$. Note that $S_{i}-S_{j}=\emptyset$ for $S\in\mathcal{C}_{1}$. For $S\in\mathcal{C}_{2}\cup\mathcal{C}_{3}$, denote by $H_{S}^{(i,j)}=\mathsf{H}(S_{i}-S_{j})$ for notational convenience. By plugging $A=S_{i}$ and $B=-S_{j}$ into (7), we have
$|S_{i}-S_{j}|\geq|S_{i}|+|S_{j}|-|H_{S}^{(i,j)}|.$ (69)
For two distinct channel indices $i,j\in\mathcal{I}_{M}$, let
$\mathcal{E}^{(i,j)}\triangleq\{S\in\mathcal{C}_{2}\cup\mathcal{C}_{3}:\,\{S_{i},S_{j}\}\text{ is an exceptional pair}\}$
denote the collections of all exceptional pairs associated with the $i$- and $j$-th channels. Note that $|S_{i}-S_{j}|\geq|S_{i}|+|S_{j}|-1$ when $\{S_{i},S_{j}\}$ is not an exceptional pair. By (69) and the disjoint-difference-array property, we have
$L$ $=|\mathbb{Z}_{L}|\geq\sum_{S\in\mathcal{C}_{2}\cup\mathcal{C}_{3}}|S_{i}-S_{j}|$
$\geq\sum_{S\in\mathcal{C}_{2}\cup\mathcal{C}_{3}}\left(|S_{i}|+|S_{j}|-1\right)-\sum_{S\in\mathcal{E}^{(i,j)}}\left(|H_{S}^{(i,j)}|-1\right).$
By the same arguments used to obtain (18), we can sum the above inequality over all distinct indices $i,j\in\mathcal{I}_{M}$ to derive
$\begin{split}M(M-1)L\geq&\ (M-1)(2w-M)|\mathcal{C}_{2}|+\sum_{S\in\mathcal{C}_{3}}(e_{S}-1)(2w-e_{S})\\ &-\sum_{i,j\in\mathcal{I}_{M},i\neq j}\sum_{S\in\mathcal{E}^{(i,j)}}\left(|H_{S}^{(i,j)}|-1\right).\end{split}$ (70)
By multiplying the inequality in (68) by (2wM)/2(2w-M)/2 and combining the result with (70), we get
$(w-1)(2w-M)(|\mathcal{C}_{1}|+|\mathcal{C}_{2}|)+\sum_{S\in\mathcal{C}_{3}}(e_{S}-1)(2w-e_{S})$
$\leq$ $\ M(M-1)L+\sum_{i,j\in\mathcal{I}_{M},i\neq j}\sum_{S\in\mathcal{E}^{(i,j)}}\left(|H_{S}^{(i,j)}|-1\right)+\frac{M(L-1)(2w-M)}{2}+(2w-M)(w-M)$
Note that we have dealt with the summation $\sum_{S\in\mathcal{C}_{3}}(e_{S}-1)(2w-e_{S})$ in the proof of Theorem 4. By (21), the above inequality can be simplified to
$|\mathcal{C}|\leq\frac{M(M-1)L+\sum_{i,j\in\mathcal{I}_{M},i\neq j}\sum_{S\in\mathcal{E}^{(i,j)}}\left(|H_{S}^{(i,j)}|-1\right)}{(2w-M)(w-1)}+\frac{M(L-1)+2(w-M)}{2w-2}.$ (71)
For $S\in\mathcal{E}^{(i,j)}$, by Lemma 2(i), $|H_{S}^{(i,j)}|\geq 2$. It follows from Proposition 1(iii) that $|H_{S}^{(i,j)}|$ divides $|S_{i}-S_{j}|$. Note that $|H_{S}^{(i,j)}|$ divides $L$ since $H_{S}^{(i,j)}$ is a subgroup of $\mathbb{Z}_{L}$. For a given subset $H$ of $\mathbb{Z}_{L}$, define $\delta(H)$ as the collection of elements $S\in\mathcal{E}^{(i,j)}$ such that $H_{S}^{(i,j)}=H$. Since, for two distinct $S,S^{\prime}\in\delta(H)$, $(S_{i}-S_{j})\cap(S^{\prime}_{i}-S^{\prime}_{j})=\emptyset$ due to the disjoint-difference-array property, we have
$\sum_{S\in\delta(H)}|S_{i}-S_{j}|\leq L.$
This implies that
$|\delta(H)|\leq\frac{L}{|H|}.$
In other words, there are at most $L/|H|$ exceptional pairs $\{S_{i},S_{j}\}$ over all $\mathcal{E}^{(i,j)}$ such that their set of stabilizers is $H$. By Lemma 2(i) again, $|H_{S}^{(i,j)}|\leq|S_{i}|+|S_{j}|-2\leq w-2$, which is strictly less than any prime factor of $L^{\prime}$. By Lagrange’s theorem, the value $|H_{S}^{(i,j)}|$ divides $(2\frac{w}{M}-1)$. Moreover, $|H_{S}^{(i,j)}|\geq 2$. That means, there are at most $\tau(2\frac{w}{M}-1)-1$ candidates for the set of stabilizers of an exceptional pair associated with $\mathcal{E}^{(i,j)}$. Hence, we conclude that
$\sum_{S\in\mathcal{E}^{(i,j)}}\left(|H_{S}^{(i,j)}|-1\right)$ $\leq\left(\tau(2\frac{w}{M}-1)-1\right)\cdot\frac{L}{|H|}\left(|H|-1\right)$
$\leq\left(\tau(2\frac{w}{M}-1)-1\right)(L-L^{\prime}),$
where the second inequality is due to $|H|$ being a factor of $(2\frac{w}{M}-1)$. Finally, we get
$\sum_{i,j\in\mathcal{I}_{M},i\neq j}\sum_{S\in\mathcal{E}^{(i,j)}}\left(|H_{S}^{(i,j)}|-1\right)\leq M(M-1)\left(\tau(2\frac{w}{M}-1)-1\right)(L-L^{\prime}).$
The result follows by (71). ∎
VII Two Extensions
VII-A At Most One-Packet Per Time Slot MC-CACs
A code $\mathcal{C}\in\text{MC-CAC}(M,L,w)$ is said to have at most one-packet per time slot (AM-OPPTS) if, for all codeword $S\in\mathcal{C}$, the element “one” appears at most once in each column [4]. That is, $|\{m:\,(m,t)\in S\}|\leq 1$ for each $t\in\mathbb{Z}_{L}$. Obviously, the AM-OPPTS property is meaningless if $L<w$. Let AM-OPPTS MC-CAC(M,L,w)(M,L,w) be the collection of codes in $\text{MC-CAC}(M,L,w)$ with AM-OPPTS property. The maximum size of a code in AM-OPPTS $\text{MC-CAC}(M,L,w)$ is denoted by $A(M,L,w)$, i.e.,
$A(M,L,w)\triangleq\max\{|\mathcal{C}|:\,\mathcal{C}\in\text{AM-OPPTS }\text{MC-CAC}(M,L,w)\}.$
By definitions, we have $A(M,L,w)\leq K(M,L,w)$ for any integers $M,L$ and $w$.
The following upper bound of $A(M,L,w)$ is due to [4, Corollaries 3.6, 3.7].
Theorem 16 ([4]).
Let $M,L$ and $w$ be positive integers. If the prime factors of $L$ are all larger than or equal to 2w12w-1, we have
$A(M,L,w)\leq\frac{M(M-1)(L-1)}{w(w-1)}+\frac{M(L-1)}{2w-2}.$ (72)
The upper bounds in Theorem 16 can be improved in the case where $M<w$, which is shown in the following theorem.
Theorem 17.
Let $M,L$ and $w$ be positive integers with $M<w$. If the prime factors of $L$ are all larger than or equal to 2w12w-1, we have
$A(M,L,w)\leq\frac{M(M-1)(L-1)}{(2w-M)(w-1)}+\frac{M(L-1)}{2w-2}.$ (73)
Proof.
The characteristic of a code $\mathcal{C}$ in AM-OPPTS $\text{MC-CAC}(M,L,w)$ is that, $0\notin D_{S}(i,j)$ for any codeword $S$ involved and any distinct indices $i,j\in\mathcal{I}_{M}$. Therefore, the inequality in (73) can be obtained in the same arguments given in the proof of Theorem 4, but just modifying (17) to be
$\sum_{S\in\mathcal{C}_{2}\cup\mathcal{C}_{3}}|D_{S}(i,j)|\leq|\mathbb{Z}^{*}_{L}|=L-1,$
which changes the last term M(M1)LM(M-1)L on the right-hand-side in (19) to M(M1)(L1)M(M-1)(L-1). Hence, the result follows. ∎
We remark that an alternate proof of Theorem 16 can be obtained by means of the arguments in Appendix A.
Based on the constructions developed in Sections V and VI, we obtain the following results.
Theorem 18.
Let $p_{1}<\cdots<p_{n}$ be primes such that $p_{i}-1$ is divisible by 2w22w-2 for each $i$. For each $1\leq i\leq n$, if there is a code in $\mathrm{CAC^{e}}(p_{i},w)$ with $(p_{i}-1)/(2w-2)$ codewords and the two conditions in (26) and(27) hold, then for any positive integers $r_{1},\ldots,r_{n}$,
$\prod_{i=1}^{n}p_{i}^{r_{i}}+\frac{\left(\prod_{i=1}^{n}p_{i}^{r_{i}}\right)-1}{w-1}-1\leq A(2,L,w)\leq\prod_{i=1}^{n}p_{i}^{r_{i}}+\frac{\left(\prod_{i=1}^{n}p_{i}^{r_{i}}\right)-1}{w-1},$
where $L=(w-1)p_{1}^{r_{i}}\cdots p_{n}^{r_{n}}$.
Proof.
The lower bound can be derived from the construction in the proof of Theorem 11 by ignoring the last codeword, say the unique one in $\mathcal{C}_{3}$. The upper bound is a consequence of Theorem 13 because $A(2,L,w)\leq K(2,L,w)$. ∎
Theorem 19.
Let $M,L,L^{\prime}$, and $w$ be positive integers such that $w>M\geq 3$, $M$ divides $w$, $L=(2\frac{w}{M}-1)L^{\prime}$, and all prime factors of $L^{\prime}$ are larger than or equal to 2w12w-1. If there is a code in $\mathrm{CAC^{e}}(L^{\prime},w)$ with $(L^{\prime}-1)/(2w-2)$ codewords, then
$\frac{M(M-1)L}{(2w-M)(w-1)}+$ $\frac{M(L-1)}{2w-2}\leq A(M,L,w)$
$\leq\frac{M(M-1)(L+(\tau(2\frac{w}{M}-1)-1)(L-L^{\prime}))}{(2w-M)(w-1)}+\frac{M(L-1)+2(w-M)}{2w-2}.$
Proof.
In the proof of Theorem 14, it is easy to see from (66) that $0\notin D_{S_{a}}(i,j)$ for $i\neq j$. This implies that the code obtained in Theorem 14 is has the AM-OPPTS property. The upper bound is a consequence of Theorem 15 because $A(M,L,w)\leq K(M,L,w)$. ∎
VII-B Mixed-Weight Codes
[5] introduced mixed-weight CACs for the purpose of increasing the throughput and reducing the access delay of some potential users with higher priority. This section will provide some constructions of mixed-weight CACs, which can be seen as generalizations of Theorem 6 and Theorem 9. Furthermore, the concept of mixed-weight will also be extended to a multichannel version by relaxing the identical-weight constraint in prior studies of MC-CACs.
Definition 5.
Let $L$ be a positive integer and $\mathcal{W}$ be a set of positive integers. A mixed-weight CAC $\mathcal{C}$ of length $L$ with weight-set $\mathcal{W}$ is a collection of subsets of $\mathbb{Z}_{L}$ such that (i) each subset is of size in $\mathcal{W}$; and (ii) $\mathcal{C}$ satisfies the disjoint-difference-set property as shown in (1).
We can use the method given in [5, Theorem 16] to construct a mixed-weight CAC of length $(w-1)p_{1}^{r_{1}}\cdots p_{n}^{r_{n}}$ with weight-set $\mathcal{W}=\{w-1,w,w_{i_{1}}^{*},w_{i_{2}}^{*},\ldots,w_{i_{t}}^{*}\}$ provided a code in $\mathrm{CAC^{e}}(p_{i_{j}}^{r_{i_{j}}},w_{i_{j}}^{*})$ exists for some $i_{1},i_{2},\ldots,i_{k}$.
Theorem 20.
Let $w,r_{1},r_{2},\ldots,r_{n}$ be positive integers and $p_{1}<p_{2}<\cdots<p_{n}$ be primes such that $p_{1}\geq w$. Suppose, for each $1\leq i\leq n$, the two conditions given in (26) and (27) hold and there exists a code $\mathcal{A}_{i}\in\mathrm{CAC}(p_{i}^{r_{i}},w_{i}^{*})$ that contains $m_{i}$ equi-difference codewords, where $w_{i}^{*}$ is an arbitrary positive integer and $m_{i}$ is not necessarily non-zero. Then there exists a code $\mathcal{C}\in\mathrm{CAC}(L,\mathcal{W})$ with
$|\mathcal{C}|=\frac{p_{1}^{r_{1}}\cdots p_{n}^{r_{n}}-1}{2}+1+\sum_{i=1}^{n}m_{i},$
where $L=(w-1)p_{1}^{r_{1}}\cdots p_{n}^{r_{n}}$ and $\mathcal{W}=\{w-1,w,w_{1}^{*},w_{2}^{*},\ldots,w_{n}^{*}\}$. In particular, if all codewords in each $\mathcal{A}_{i}$ are non-exceptional, then $\mathcal{C}$ contains $1+\sum_{i=1}^{n}m_{i}(w_{i}^{*}-1)$ codewords with weight w1w-1, $\frac{p_{1}^{r_{1}}\cdots p_{n}^{r_{n}}}{2}-\sum_{i=1}^{n}m_{i}(w_{i}^{*}-1)$ codewords with weight $w$, and $m_{i}$ codewords with weight $w_{i}^{*}$ for each $i$.
Theorem 21.
Let $p_{1}<\cdots<p_{n}$ be primes and $w_{1},\ldots,w_{n}$ be positive integers (not necessarily distinct) such that $p_{i}\geq 2w_{i}-1$ for any $i$. If there is a code in $\mathrm{CAC^{e}}(p_{i},w_{i})$ with $m_{i}$ codewords for each $i$, then for any positive integers $r_{1},\ldots,r_{n}$, there exists a mixed-weight code $\mathcal{C}\in\mathrm{CAC}(L,\mathcal{W})$ with
$|\mathcal{C}|=\sum_{i=1}^{n}\frac{m_{i}(p_{i}^{r_{i}}-1)}{p_{i}-1}\prod_{j=i+1}^{n}p_{j}^{r_{j}},$
where $L=p_{1}^{r_{1}}\cdots p_{n}^{r_{n}}$ and $\mathcal{W}=\{w_{1},\ldots,w_{n}\}$. In particular, $\mathcal{C}$ contains
$\frac{m_{i}(p_{i}^{r_{i}}-1)}{p_{i}-1}\prod_{j=i+1}^{n}p_{j}^{r_{j}}$
codewords with weight $w_{i}$, for $i=1,2,\ldots,n$.
Proof.
The proof can be done with the same arguments in the proof of Theorem 9 by adjusting $\Gamma_{i}$ being the set of $m_{i}$ generators of the given code in $\mathrm{CAC^{e}}(p_{i},w_{i})$ and replacing the definition of $S_{a}$ in (37) with
$S_{a}=\{ja\in\mathbb{Z}_{p_{1}^{r_{1}}}\times\cdots\times\mathbb{Z}_{p_{n}^{r_{n}}}:\,j=0,1,2,\ldots,w_{i}-1\}.$
We now extend the concept of mixed-weight to multichannel CACs.
Definition 6.
Let $M,L$ be two positive integers and $\mathcal{W}$ be a set of positive integers. A mixed-weight MC-CAC $\mathcal{C}$ of length $L$ with weight-set $\mathcal{W}$ in $M$ channels is a collection of subsets of $\mathcal{I}_{M}\times\mathbb{Z}_{L}$ such that (i) each subset is of size in $\mathcal{W}$; and (ii) $\mathcal{C}$ satisfies the disjoint-difference-array property as shown in (3).
VIII Concluding Remarks
It has been shown in [1, Theorem 1] that
$\limsup_{L\to\infty}\frac{K(L,w)}{L}=\frac{1}{2w-2}.$ (74)
When considering multichannel CACs, if $M=2$, it reveals from Theorem 13 that
$\frac{K(2,L,w)}{L}\sim\frac{1}{w-1}+\frac{1}{(w-1)^{2}}\quad\text{as}\quad L\to\infty$ (75)
under some specific parameter settings. For general $M\geq 3$, when $w>M$ and under other specific parameter settings, we can see from Theorem 14 that the lower bound of $K(M,L,w)/L$ approaches
$\frac{M(M-1)}{(2w-M)(w-1)}+\frac{M}{2w-2}$ (76)
as $L$ goes to infinity. However, based on Theorem 15, our derived upper bound of $K(M,L,w)/L$ approaches
$\frac{M(M-1)\left(1+(\tau(m)-1)(1-\frac{1}{m})\right)}{(2w-M)(w-1)}+\frac{M}{2w-2}$ (77)
as $L$ goes to infinity, where $m=2\frac{w}{M}-1$. One can see that the asymptotic value in (76) reduces to (74) and (75) when $M=1$ and $2$, respectively. Motivated by (74), we have the following conjecture.
Conjecture 1.
For all $w\geq M$, we have
$\limsup_{L\to\infty}\frac{K(M,L,w)}{L}=\frac{M(M-1)}{(2w-M)(w-1)}+\frac{M}{2w-2}.$
This conjecture remains an open problem for future investigation.
Appendix A A New Proof of Theorem 3
Let $\mathcal{C}$ be any code in $\text{MC-CAC}(M,L,w)$. Similar to the proof of Theorem 4, all codewords are non-exceptional and partitioned into the three classes
$\mathcal{C}_{1}$ $=\{S\in\mathcal{C}:e_{S}=1\},$
$\mathcal{C}_{2}$ $=\{S\in\mathcal{C}:e_{S}=w\},\text{ and}$
$\mathcal{C}_{3}$ $=\{S\in\mathcal{C}:2\leq e_{S}\leq w-1\}.$
By the same argument in the proof of Theorem 4, we have
$(w-1)|\mathcal{C}_{1}|+\sum_{S\in\mathcal{C}_{3}}(w-e_{S})\leq\frac{M(L-1)}{2},$ (78)
and
$w(w-1)|\mathcal{C}_{2}|+\sum_{S\in\mathcal{C}_{3}}(e_{S}-1)(2w-e_{S})\leq M(M-1)L,$ (79)
which are analogous to (16) and (18), respectively. Multiplying the inequality in (78) by $w$ and combining the result with (79), we obtain
$w(w-1)(|\mathcal{C}_{1}|+|\mathcal{C}_{2}|)+\sum_{S\in\mathcal{C}_{3}}\left(-e_{S}^{2}+(w+1)e_{S}+w^{2}-2w\right)\leq\frac{wM(L-1)}{2}+M(M-1)L.$ (80)
Define a function $f(x)=-x^{2}+(w+1)x+w^{2}-2w$ on the interval [2,w1][2,w-1]. It is not hard to see $f(x)$ has the minimum value $w^{2}-2$ at $x=2$ or x=w1x=w-1, which leads to
$f(x)\geq w^{2}-2\geq w(w-1)$ (81)
because of $w\geq 2$. Hence, it follows from (80) that
$w(w-1)|\mathcal{C}|\leq\frac{wM(L-1)}{2}+M(M-1)L,$ (82)
as desired.
Note that in (81), f(x)>w(w1)f(x)>w(w-1) if $w\geq 3$, leading to a strict inequality in (82). This concludes that the equality (14) holds only when all involved codewords are either in $\mathcal{C}_{1}$ or in $\mathcal{C}_{2}$, i.e., the packets must either be on the same channel, or else they have to be distributed across different channels.